### 16 Mar 2026 v1

# Minimum-Action Learning: Energy-Constrained Symbolic Model Selection for Physical Law Identification from Noisy Data

Abstract
Identifying physical laws from noisy observational data is a central challenge in scientific machine learning. We present Minimum-Action Learning (MAL), a framework that selects symbolic force laws from a pre-specified basis library by minimizing a Triple-Action functional combining trajectory reconstruction ($I_{\mathrm{max}}$), architectural sparsity ($E_{\mathrm{min}}$), and energy-conservation enforcement ($\mathcal{L}_{\mathrm{Symmetry}}$). A wide-stencil acceleration-matching technique reduces noise variance by $10^{4}\times$, transforming an intractable problem (SNR $\approx 0.02$) into a learnable one (SNR $\approx 1.6$); this preprocessing is the critical enabler shared by all methods tested, including SINDy variants. Using two benchmarks—Kepler (inverse-square gravity) and Hooke’s law (linear restoring force)—we show that MAL recovers the correct force law with Kepler exponent $p=3.01\pm 0.01$ in 835 s at $\sim$0.07 kWh (40% reduction vs. prediction-error-only baselines). The raw correct-basis rate is 40% for Kepler and 90% for Hooke; an energy-conservation-based model selection criterion discriminates the true force law from alternatives in all cases, yielding 100% pipeline-level identification. A physics-informed gate initialization independently achieves 10/10 correct selection. Direct comparison against noise-robust SINDy variants, HNNs, and LNNs confirms MAL’s distinct niche: energy-constrained, interpretable model selection combining symbolic basis identification with dynamical rollout validation. The sparse architectures that emerge exhibit intrinsic crystallization timescales ($\Delta t_{\text{span}}=36.2\pm 4.1$ epochs) independent of initialization. The regularization design draws structural inspiration from biological metabolic constraints, though the biological analogies discussed remain suggestive rather than formally established.
Keywords: neural architecture search, action principles, energy optimization, physical law identification, modularity, developmental neuroscience
The principle of least action, formulated by Maupertuis and refined by Lagrange and Hamilton, has unified physics from classical mechanics to quantum field theory [14]. Simultaneously, biological systems—constrained by thermodynamics and natural selection—have evolved architectures that minimize metabolic expenditure while maximizing information processing [44, 25]. We propose that these parallel optimization strategies are not coincidental but reflect a deeper organizing principle operating across scales: network-weighted action minimization. Here we demonstrate that embedding this principle in artificial neural networks—via a Triple-Action functional combining information maximization ($I_{\mathrm{max}}$), energy minimization ($E_{\mathrm{min}}$), and symmetry enforcement—enables energy-constrained identification of physical laws from noisy data, while spontaneously generating the modular architectures characteristic of evolved systems.
Introduction
The Energy-Physics Connection
From Claude Bernard’s milieu intérieur [2] through Cannon’s homeostasis [5] to contemporary network physiology [1], biology has recognized that living systems operate through energy-constrained multi-scale integration. Kleiber’s discovery [25] that metabolic rate scales as $M^{3/4}$ across 17 orders of magnitude in body mass, later explained by West, Brown, and Enquist [44] through network optimization, demonstrates that metabolic constraints impose mathematical regularities on biological organization.
In parallel, physics has long relied on action principles—the requirement that systems follow paths extremizing $S=\int L\,dt$, where $L$ is the Lagrangian—to derive equations of motion [17]. From classical mechanics to quantum field theory, nature appears to solve optimization problems, minimizing action subject to symmetry constraints [43].
Recent work in developmental neuroscience reveals that these principles converge in the brain: glial cells enforce metabolic constraints on neural architecture search during development [19, 42], while dynamic coordination theory [34, 22, 23, 24] formalizes the relationship between energy states and network synchronization through metastable dynamics [38] and Farey-tree patterns—a framework supported by nonlinear frequency-locking phenomena in autonomic neural circuits [16] and phase-oscillator models of synaptic plasticity [6]. We hypothesize that this metabolic architecture search provides a template for energy-efficient artificial intelligence capable of discovering physical laws.
The Kepler Benchmark: Rediscovering $F=Gm_{1}m_{2}/r^{2}$
Newton’s derivation of the inverse-square law from Kepler’s observations represents one of science’s foundational achievements. Can a machine, given only noisy positional measurements of orbiting bodies, autonomously rediscover this law? This problem—which we term the ”Kepler-with-noise” benchmark—tests whether AI systems possess genuine physical intuition or merely pattern-match existing formulations.
Existing approaches address aspects of this challenge. Symbolic regression methods [40, 11] and GNN-based symbolic distillation [11, 26] can recover force laws from clean or moderately noisy data, but degrade at high noise levels. Hamiltonian Neural Networks (HNNs) [18] and Lagrangian Neural Networks (LNNs) [10] embed conservation structure but learn black-box energy functions rather than selecting interpretable symbolic forms. Physics-informed neural networks [32] require pre-specifying the equations to solve. Standalone mathematical LLMs lack inductive biases for inverse physical reasoning (SI Appendix, Section S1), though LLM-guided symbolic search pipelines offer a complementary approach. Noether’s Razor [41] and Noether’s Learning Dynamics [37] provide complementary theoretical perspectives on symmetry-driven model selection.
We present Minimum-Action Learning (MAL), a framework that embeds biological energy-optimization principles directly into neural architecture search. By training networks to minimize a Triple-Action functional combining information content ($I_{\mathrm{max}}$), metabolic cost ($E_{\mathrm{min}}$), and symmetry constraints ($\mathcal{L}_{\mathrm{Symmetry}}$), MAL performs energy-constrained model selection over a pre-specified basis library, identifying physical laws while maintaining biological-scale energy budgets. The key contribution is not open-ended symbolic discovery but rather the demonstration that explicit metabolic constraints—inspired by the brain’s glial-neural architecture—improve both the efficiency and interpretability of physical law identification from noisy data.
Results
Energy-Constrained Identification of Newton’s Law
We implemented MAL as a differentiable neural network (MinActionNet) incorporating three key architectural innovations (Fig. 2, Methods):
(1) Noether Force Basis. To enforce rotational symmetry (SO(2) invariance), we parameterize forces as radial functions: $\mathbf{F}(\mathbf{r})=f(r)\hat{\mathbf{r}}$, where $f(r)=\sum_{i=1}^{5}A_{i}\theta_{i}\phi_{i}(r)$ and $\phi_{i}\in\{r^{-2},r^{-1},r,1,r^{-3}\}$ is a library of candidate basis functions. Learnable gates $A_{i}=\mathrm{softmax}(\ell_{i}/\tau)$, where $\ell_{i}$ are gate logits, select among bases via temperature-annealed competition.
(2) Triple-Action Objective. The loss function implements three constraints:
$\mathcal{L}=\alpha_{I}\mathcal{L}_{I_{\max}}+\alpha_{E}\mathcal{L}_{E_{\min}}+\alpha_{S}\mathcal{L}_{\mathrm{Symmetry}}$ (1)
where $\mathcal{L}_{I_{\max}}$ combines trajectory reconstruction and wide-stencil acceleration matching (maximizing information extraction from noisy data), $\mathcal{L}_{E_{\min}}$ penalizes architectural complexity via gate entropy and coefficient sparsity (minimizing energy), and $\mathcal{L}_{\mathrm{Symmetry}}$ enforces energy conservation (Noether’s theorem).
(3) Bimodal Glial-Neural Optimization (BGNO). Drawing a structural analogy from the brain’s 1:1 glia-to-neuron ratio [9] and evidence that glial cells mediate metabolic constraints during neural development [7, 12], we implement a two-phase training schedule in which the Triple-Action functional drives the architecture $\mathbf{A}(t)$ along a soft-to-discrete manifold, where energy-driven sharpening ($E_{\mathrm{min}}$) encourages the emergence of discrete structural motifs from initially uniform gate distributions. During warmup (50 epochs, low regularization $\alpha_{E}=0.01$, uniform gates $\tau=1.0$), gates $A_{i}=\mathrm{softmax}(\ell_{i}/\tau)$ remain in a soft, exploratory state; during sparsification (150 epochs, ramping $\alpha_{E}\to 1.0$, annealing $\tau\to 0.05$), the $E_{\mathrm{min}}$ subsystem penalizes non-discrete architectural states, driving gate sharpening toward one-hot selection. This mechanism is analogous to temperature-annealed differentiable architecture search [45], here motivated by biological metabolic constraints rather than computational efficiency alone. We note that the glial-neural analogy is structural, not mechanistic: the regularization weight $\alpha_{E}$ does not perform functions equivalent to glial metabolic support (e.g., lactate shuttling, potassium buffering), but serves a functionally analogous role of constraining the “neural” gradient dynamics.
Figure 1: Trajectory reconstruction and Hamiltonian conservation. (A) Comparison of ground-truth Keplerian orbits (blue) vs. MAL reconstruction (red) for a test orbit rolled out over 5 orbital periods from initial conditions using the identified $r^{-2}$ force law. Slight enlargement is attributable to the 6% deficit in recovered GMGM. (B) Energy conservation error $\Delta H$ remains bounded, enforced by the Noether-symmetry term $\mathcal{L}_{\mathrm{Symmetry}}$ in the Triple-Action functional, which constrains the parameter trajectory $\theta(t)$ to satisfy $dH/dt\approx 0$. Implementation: the force is computed via NoetherForceBasis.forward(), which evaluates $f(r)=\sum_{i}A_{i}\theta_{i}\phi_{i}(r)$ with gates $A_{i}=\mathrm{softmax}(\texttt{A\_logits}/\tau)$.
Figure 2: Soft-to-discrete architectural crystallization. (A) Evolution of gate activation probabilities $A_{i}$ from uniform (epoch 1, soft state) to one-hot selection of the $r^{-2}$ basis (epoch 200, discrete state). This transition occurs on a soft-to-discrete architecture manifold where A_logits are sharpened via softmax temperature decay ($\tau:1\to 0.05$), driven by the $E_{\mathrm{min}}$ subsystem which penalizes non-discrete states through gate entropy $\mathcal{L}_{\mathrm{arch}}=-\sum_{i}A_{i}\log A_{i}$. (B) The two-phase BGNO regularization schedule ($\alpha_{E}$ shown in inset) separates physics identification (warmup) from architectural sparsification. Shown for a representative seed (seed 0); variability across 10 seeds is reported in SI Table S1.
Training on 16 synthetic Keplerian orbits (semi-major axes $a\in[0.5,5]$ AU, eccentricities $e\in[0,0.3]$, observation noise $\sigma=1\%$ of median radius) for 200 epochs yielded:
  • Correct basis selection: Gate probability $A_{0}=1.000$ for the $r^{-2}$ basis in the primary run (Fig. 2). Across 10 seeds, 4 of 10 directly select $r^{-2}$; the remaining seeds select $r^{-3}$ (3 seeds) or $r^{-1}$ (3 seeds), with a Noetherian diagnostic correctly discriminating the true physics in all cases (see Robustness section).
  • Accurate force calibration: Recovered gravitational coupling $\theta_{0}=0.936$ (true: $GM=1.0$), representing 6.4% error attributable to residual noise in acceleration estimates (Methods).
  • Kepler’s third law verification: Power-law fit to orbital periods $T^{2}\propto a^{p}$ yielded $p=3.01\pm 0.01$ (theoretical: 3.0), confirming that the identified force law generates correct dynamics (Fig. 1).
  • Energy efficiency: Total training time 835 seconds, energy consumption $\sim$0.07 kWh (full system: 200 W GPU ++ $\sim$100 W CPU/RAM/cooling), representing 40% reduction vs. prediction-error-only baselines (SI Appendix, Table S2).
[…]
The Noise Bottleneck and Wide-Stencil Solution
A critical insight emerged from failure analysis: naive finite-difference acceleration estimates from noisy positions have signal-to-noise ratio (SNR) $\ll 1$, rendering learning impossible. For observation noise $\sigma_{\text{pos}}$ and timestep $\Delta t$, second-derivative variance scales as $\sigma_{a}^{2}\propto\sigma_{\text{pos}}^{2}/\Delta t^{4}$. At typical scales ($\sigma_{\text{pos}}=0.016$ AU, $\Delta t=0.05$), this yields $\sigma_{a}\approx 15.7$, while true gravitational acceleration at $r=2$ AU is $\approx 0.25$—an SNR of $0.016$ (Table 1, SI Appendix).
We solved this through wide-stencil acceleration matching: computing second derivatives with stride $s=10$ reduces noise variance by $s^{4}=10{,}000\times$ while introducing only $O(s^{2}\Delta t^{2})$ systematic error ($0.3\%$ for Keplerian orbits). This transforms an intractable problem (SNR $\approx 0.02$) into a learnable one (SNR $\approx 1.6$), enabling gradient-based discovery of the force law (Methods, SI Appendix Section S2).
Table 1: Noise reduction via wide-stencil differentiation
Stride $s$ Noise $\sigma_{a}$ Signal $|a|$ SNR Training
1 (naive) 15.7 0.25 0.016 Failed
5 0.63 0.25 0.40 Partial
10 0.16 0.25 1.6 Success
20 0.039 0.25 6.4 Success
Robustness, Success Rate, and Intrinsic Timescales
To assess reliability, we trained 10 independent models with different random seeds on identical data. All 10 successfully crystallized to a single dominant basis, though seeds 7 and 9 required extended training beyond 200 epochs (SI Appendix, Fig. S1–S4, Table S1).
Correct-basis success rate. Of 10 seeds, 4 directly selected the correct $r^{-2}$ basis (seeds 0, 1, 2, 8); 3 selected $r^{-3}$ (seeds 3, 4, 5) and 3 selected $r^{-1}$ (seeds 6, 7, 9). The raw correct-basis rate is thus 40%. However, all 10 models recovered Kepler exponent $p\in[2.995,3.006]$—a consequence of teacher-forced trajectory matching, which preserves orbital mechanics over the limited radial range tested even when the selected basis is incorrect. This means $p\approx 3.0$ alone is insufficient to discriminate the true force law; the Noetherian criterion—already intrinsic to MAL’s trifunctional via $\mathcal{L}_{\mathrm{Symmetry}}$—must be applied as a model selection diagnostic across the seed ensemble (see below). Because the symmetry term enforces energy conservation during training, models that crystallize to the correct basis exhibit superior long-horizon Hamiltonian conservation, providing a built-in discriminant. Applying this criterion across seeds achieves 100% identification of the correct physics (SI Table S2).
Basis selection interventions. To diagnose the 40% direct selection rate, we tested three interventions across 10 seeds each (SI Table S4): (i) extended warmup (100 epochs, 4/10 correct), (ii) reduced noise ($\sigma=0.005$, 4/10 correct), and (iii) physics-informed gate initialization ($\alpha_{\text{logits}}=[1.5,0,0,0,0]$, giving $r^{-2}$ approximately 50% initial gate weight). Neither longer exploration nor cleaner data improved the success rate, but biased initialization achieved 10/10 correct selection, confirming that gate competition during the warmup phase—not noise level or exploration time—is the bottleneck. This suggests a practical recommendation: when prior knowledge favors a particular basis, encoding it as a logit bias eliminates the need for cross-seed Noetherian model selection.
Intrinsic timescales. Three invariant quantities emerged across seeds:
(1) Crystallization span: Once gate selectivity $R=A_{\max}/A_{\text{2nd}}$ exceeds 10 (onset), it reaches $R>1000$ (frozen) after $\Delta t_{\text{span}}=36.2\pm 4.1$ epochs (bootstrap 95% CI: [32.8, 39.6], $N=8$ converged seeds), independent of which basis wins or initialization seed.
(2) Growth rate: Within the crystallization window, $R$ grows geometrically at rate $\gamma=1.137\pm 0.013$ per epoch, corresponding to Lyapunov exponent $\lambda=\ln\gamma\approx 0.128$ epoch-1.
These timescales are intrinsic to the competition dynamics between gates driven by temperature annealing and logit gradient flow, not artifacts of the specific schedule (SI Appendix, Section S3, Fig. S5).
Generalization to Hooke’s Law
To test whether MAL generalizes beyond inverse-square gravity, we applied it to Hooke’s law (F=krF=-kr, linear restoring force) using the same basis library $\{r^{-2},r^{-1},r,1,r^{-3}\}$, identical architecture, and identical training protocol (SI Appendix, Section S4A, SI Table S3). Of 10 seeds, 9 directly selected the correct $r$ basis (90% vs. 40% for Kepler), recovering spring constant $\hat{k}=0.980\pm 0.001$ (true: $k=1.0$, 2% error). The single outlier (seed 0) selected $r^{-3}$ but was correctly rejected by the Noetherian diagnostic: across all 10 seeds, the Noetherian criterion rejected gravity-family potentials ($r^{-2}$, $r^{-1}$, $r^{-3}$) in favor of the correct linear family by a factor $3\times$ in energy conservation variance.
The higher direct-selection success rate (90% vs. 40%) reflects Hooke’s simpler competition landscape—the loss-surface topography over gate logit space that determines which basis functions attract gradient flow during warmup. For Hooke, the linear basis $r$ is the only function in the library with the correct radial scaling at the near-circular orbits used, creating a single dominant attractor; for Kepler, $r^{-2}$ and $r^{-3}$ are near-degenerate over limited radial ranges, creating competing attractors that trap 6 of 10 seeds. Crystallization timescales ($\Delta t_{\text{span}}=42.5\pm 1.6$ epochs, $\gamma=1.113\pm 0.003$) are comparable to Kepler ($36.2\pm 4.1$ epochs), supporting the universality of the gate competition dynamics.
Table 2: Summary of MAL pipeline performance across benchmarks
Kepler ($r^{-2}$) Hooke ($r$)
Basis library $\{r^{-2},r^{-1},r,1,r^{-3}\}$
Direct selection rate 4/10 (40%) 9/10 (90%)
Pipeline rate 10/10 (100%) 10/10 (100%)
Biased init rate 10/10 (100%)
Recovered coeff. $\hat{\theta}_{0}=0.94$ $\hat{k}=0.98$
Kepler exponent $\hat{p}$ $3.01\pm 0.01$
Training energy 0.07 kWh 0.07 kWh
Time/seed 835 s $\sim$835 s
After energy-conservation-based cross-seed selection.
Energy-Conservation-Based Model Selection
Because MAL’s trifunctional includes an energy-conservation term $\mathcal{L}_{\mathrm{Symmetry}}=\mathrm{Var}[E]$ (inspired by Noether’s theorem linking time-translation symmetry to energy conservation), models that crystallize to the correct basis are trained to conserve energy—providing a built-in discriminant that can be applied as a model selection criterion across seeds without any additional computation. For long-horizon rollouts (5 orbital periods), we computed Hamiltonian variance $\sigma_{H}^{2}=\langle(H-\langle H\rangle)^{2}\rangle$: models selecting the correct $r^{-2}$ basis conserve energy $3\times$ better than $r^{-1}$ and $6\times$ better than $r^{-3}$ models, despite all achieving comparable short-term trajectory accuracy (SI Table S2, Fig. 4).
Figure 4: Energy-conservation-based model selection and schedule geometry. (Left) Phase-space trajectory of $(\alpha_{E},\tau)$ during training, color-coded by epoch. Red diamonds mark epochs where the ratio $\alpha_{E}/\tau$ passes through integer values (3:1, 2:1, 1:1), coinciding with major transitions in gate selectivity (onset, sparsification, crystallization). These coincidences arise from the designed schedule geometry; whether they reflect deeper dynamical principles remains an open question (see Discussion). (Right) Energy conservation $\sigma_{H}$ by selected basis: correct $r^{-2}$ models conserve $H$ over long-horizon rollouts, while incorrect bases violate Hamiltonian dynamics despite matching short-term trajectories. This provides an energy-conservation-based diagnostic for model selection.
Schedule geometry and gate transitions. Analysis of the designed schedule trajectory $(\alpha_{E}(t),\tau(t))$ reveals that major gate transitions coincide with epochs where the ratio $\alpha_{E}/\tau$ passes through specific integer and near-integer values: gate onset ($R\geq 10$) near $\alpha_{E}/\tau\approx 2\text{:}3$, rapid sparsification near $\alpha_{E}/\tau\approx 1\text{:}1$, and final crystallization ($R\geq 1000$) near $\alpha_{E}/\tau\approx 2\text{:}1$ (Fig. 4). We emphasize that these coincidences arise from the designed schedule geometry: because $\alpha_{E}$ ramps linearly and $\tau$ decays exponentially, any smooth trajectory through this parameter space will necessarily pass through integer-ratio nodes (a consequence of the density of rationals). Whether the gate dynamics are genuinely sensitive to these passages—as opposed to merely coincident with them—remains an open question requiring perturbation experiments (see Discussion).
Architectural sparsification. The $E_{\mathrm{min}}$-driven soft-to-discrete transition produces near-one-hot gate distributions: gate concentration $C_{\mathrm{gate}}=0.99\pm 0.02$ (Herfindahl–Hirschman Index on effective gate contributions $A_{i}|\theta_{i}|$; $C_{\mathrm{gate}}=1$ indicates complete concentration on a single basis) for MAL-trained models ($N=10$ seeds) vs. $C_{\mathrm{gate}}=0.14\pm 0.04$ for teacher-forcing-only baselines (Mann–Whitney $U=100$, $p<10^{-4}$; permutation test $p<0.001$, n=10,000n=10{,}000; SI Appendix, Fig. S6). This architectural sparsification is consistent with Clune et al.’s [8] finding that minimizing wiring costs—a proxy for metabolic expenditure—produces modular structure in evolved networks.
Structural parallels across domains. We note intriguing structural parallels between MAL’s modularity emergence and patterns in biological systems, while emphasizing that these are currently analogies rather than formal equivalences:
(i) Perinatal brain development: Neonatal heart rate and breathing movement coordination during quiet (low-energy) vs. active (high-energy) sleep states exhibit integer-ratio (3:1) phase relationships [21, 33], reflecting dynamic coordination of coupled physiological oscillators under energy constraints. This frequency-locking behavior is a manifestation of the metastable coordination dynamics described by Kelso [22, 38, 24], and parallels the nonlinear frequency-locking of baroreceptor and sympathetic rhythms observed in autonomic neural circuits [16].
(ii) Ecological networks: Microbial gene co-expression networks in the TARA Oceans dataset [36] exhibit modular architectures consistent with wiring-cost minimization [8], though the connection to MAL’s training dynamics remains analogical (SI Appendix, Section S6).
We emphasize that these cross-domain parallels are structural analogies, not established causal connections. The training schedule parameters $(\alpha_{E},\tau)$ are externally imposed hyperparameters, not emergent oscillators; and the biological systems involve genuine coupled-oscillator dynamics. Whether energy-constrained optimization generically produces similar architectural motifs across these domains is a hypothesis warranting formal investigation.
Comparison to Alternative Approaches
We benchmarked MAL against five alternatives on the Kepler-with-noise problem (SI Appendix, Section S4, SI Tables S5–S6):
(1) SINDy variants [3, 20, 13]: Vanilla SINDy with naive differentiation (stride $s=1$) fails completely (noise-dominated). However, when equipped with the same wide-stencil preprocessing ($s=10$) used in MAL, vanilla SINDy identifies $r^{-2}$ in 10/10 seeds, GP-SINDy in 8/10, and ensemble-SINDy in 10/10 (SI Table S5). The wide-stencil technique is the critical enabler, not the learning algorithm. SINDy’s advantage is speed (<1<1 s vs. $\sim$835 s); its limitation is the absence of dynamical validation—SINDy returns sparse coefficients but cannot roll out trajectories or verify energy conservation. MAL’s Noetherian diagnostic provides this additional layer.
(2) HNN [18] and LNN [10]: We trained both on identical data (SI Table S6). HNN (17K parameters, 113 s) achieved excellent energy conservation ($\sigma_{H}=4.1\times 10^{-4}$) but learns a black-box Hamiltonian with no interpretable symbolic form. LNN (17K parameters, 242 s) suffered from Hessian singularities, with validation loss diverging to $10^{6}$ throughout training—a known failure mode for LNNs on noisy data where the mass matrix becomes ill-conditioned. Neither method performs basis selection or yields a symbolic force law.
(3) Mathematical LLM (Qwen2-Math 7B): 0% accuracy on inverse problems, 61% on forward derivation—mathematical manipulation without physical intuition (SI Appendix, Section S1).
(4) Physics-informed NN [32]: 85% accuracy when the inverse-square law is pre-specified, but cannot identify the law de novo.
(5) Teacher-forcing only (no $\mathcal{L}_{E_{\min}}$): Gates remained at 45% (failed to crystallize); energy consumption 77% higher.
Summary of niche. MAL occupies a distinct position: it combines interpretable symbolic basis selection (shared with SINDy [3]) with dynamical rollout and Noetherian validation (shared with HNN [18]), while adding explicit energy-efficiency optimization. HNNs/LNNs [18, 10] guarantee conservation but produce black-box energy functions; SINDy [3] yields symbolic expressions but without dynamical consistency checks; MAL provides both within a metabolically constrained framework.
Discussion
Key Findings
Our results establish four key findings:
(1) Metabolic constraints improve physical law identification. Embedding energy minimization ($E_{\min}$) alongside information maximization ($I_{\max}$) in the Triple-Action training objective constrains the model selection problem from an unconstrained search to a tractable optimization. The $E_{\mathrm{min}}$ subsystem drives gate crystallization (basis selection) that does not occur under prediction error alone, while reducing training energy by 40%. This parallels Noble’s biological relativity [28]: information content and metabolic cost jointly determine which model structure the network selects.
[…]
(3) Integer-ratio coincidences suggest cross-scale connections. We observe that major gate transitions coincide with epochs where the designed schedule ratio $\alpha_{E}/\tau$ passes through integer values. We emphasize that this observation is currently correlational: $\alpha_{E}$ and $\tau$ are externally imposed hyperparameters, not emergent oscillators, and the coincidence may reflect the density of rationals rather than dynamical phase-locking. However, the structural parallel to integer-ratio coordination in physiological oscillators [21, 16] and the broader framework of metastable coordination dynamics [34, 22, 38, 23, 24] is intriguing and motivates formal investigation. Recent work on phase-oscillator networks with synaptic plasticity [6] suggests that frequency-locking phenomena may be generic in systems where coupling weights co-evolve with oscillator dynamics—a structural analogue to MAL’s co-evolution of gate logits and force coefficients. A rigorous test would require showing that the crystallization dynamics are sensitive to the schedule trajectory’s passage through integer ratios—e.g., via circle-map analysis of the coupled $\alpha_{E}$-$\tau$ dynamics—which we leave to future work.
(4) Architectural sparsification as a consequence of energy minimization. Clune et al. [8] demonstrated that networks evolving under pressure to minimize connection costs spontaneously develop modular architectures. Our work shows a related pattern in physics model selection: MAL’s sparse gate selection ($C_{\mathrm{gate}}=0.99$) emerges from $\mathcal{L}_{E_{\min}}$, not from a pre-programmed bias. This is consistent with the hypothesis that metabolic constraints on wiring costs drive architectural sparsification across both biological [4] and artificial systems [8].
Implications for AI and Theoretical Biology
Energy-efficient AI. MAL’s 40% energy reduction over prediction-error-only baselines demonstrates that explicit metabolic constraints ($E_{\min}$ optimization) can improve computational efficiency without sacrificing task performance. The mechanism is architectural: $\mathcal{L}_{E_{\min}}$ drives early gate crystallization, reducing the effective parameter count and shortening convergence time. This principle—that sparsity-inducing regularization simultaneously improves interpretability and efficiency—may generalize beyond the current benchmarks to larger-scale scientific discovery tasks.
Energy-conservation-based model selection. A persistent challenge in AI interpretability is distinguishing genuine causal understanding from correlation [30]. The energy-conservation criterion—preferring models that preserve Hamiltonian dynamics over long-horizon rollouts—provides a physics-grounded approach complementary to recent work on learning conserved quantities [41, 37]: models that violate energy conservation reveal their failure to capture true causal structure, even when matching short-term predictions. This criterion is inspired by Noether’s theorem (time-translation symmetry implies energy conservation), though our implementation is a specific instantiation rather than a general Noether construction.
Developmental origins of health and disease. The structural analogy between MAL’s two-phase optimization and brain development suggests testable hypotheses: Does disruption of metabolic coordination (e.g., maternal inflammation [15]) alter the integer-ratio structure of physiological coupling? Could this predict neurodevelopmental outcomes? While the correspondence is currently analogical, the intrinsic crystallization timescales observed in MAL motivate quantitative investigation of equivalent timescales in biological development.
Limitations and Future Directions
Fixed basis library. MAL requires pre-specifying candidate force laws $\{\phi_{i}\}$, with the correct answer included. This makes MAL a model selection framework, not an open-ended discovery system. Basis library sensitivity experiments (SI Appendix, Section S4F) quantify this limitation: adding near-confounders ($r^{-2.5}$, $r^{-1.5}$) reduces the correct-basis rate from 50% to 20%, while removing the correct basis ($r^{-2}$) causes the system to split between alternatives. Distant additions ($r^{2}$, $r^{-4}$, $\ln r$) have no effect. In all cases, the energy-conservation diagnostic remains informative, and uniformly elevated $\sigma_{H}$ across seeds flags potential library inadequacy. Extending to genuinely autonomous discovery—where the algorithm constructs novel functional forms not in the library—requires integration with open-ended symbolic regression methods [40, 11] or LLM-guided symbolic search, which is an important direction for future work.
Two benchmarks, limited scope. We have validated MAL on two central-force problems (Kepler and Hooke) in 2D with synthetic data. The generality of our claims would be substantially strengthened by additional benchmarks: non-central forces (e.g., magnetic fields), dissipative systems (e.g., damped oscillators), higher dimensions, coupled multi-body problems, and semi-real data (e.g., planetary ephemerides with actual measurement noise). The core principle—embedding action minimization in differentiable networks—should transfer, but this remains to be demonstrated.
Correct-basis success rate. Without logit initialization, only 4 of 10 Kepler seeds directly select $r^{-2}$, necessitating cross-seed application of the energy-conservation-based diagnostic. Biased initialization resolves this (10/10), but requires prior knowledge of which basis to favor. Understanding the logit-gradient dynamics that determine gate competition—and developing initialization-free methods to improve the raw success rate—remain priorities.
Cross-scale analogies. The structural parallels between MAL’s training dynamics, physiological coordination, and ocean genomic modularity are currently analogies, not formal equivalences. The training schedule parameters are externally imposed, not emergent oscillators. Establishing formal connections would require, e.g., circle-map analysis showing sensitivity of crystallization to integer-ratio passages, or demonstrating Farey coordination in systems where the schedule is itself learned rather than designed.
Conclusion
We have demonstrated that minimum-action learning—neural network training guided by a Triple-Action functional combining information maximization, energy minimization, and symmetry enforcement—enables energy-constrained identification of physical force laws from noisy data. Validated on two benchmarks (Kepler gravity and Hooke’s linear restoring force), the key technical contributions are: (1) the wide-stencil noise reduction technique that transforms an intractable problem (SNR $\sim 0.02$) into a learnable one (SNR $\sim 1.6$); (2) the soft-to-discrete gate sharpening mechanism that achieves basis selection through energy-driven architectural crystallization, with a physics-informed initialization achieving 10/10 correct selection; and (3) the Noetherian model selection criterion that discriminates correct from incorrect physics via long-horizon energy conservation. Direct comparison against SINDy variants, HNNs, and LNNs confirms MAL’s distinct niche: interpretable, energy-constrained model selection combining symbolic basis identification with dynamical validation.
The modular architectures that emerge from this process parallel patterns observed in biological systems under metabolic constraints—including physiological coordination [21] and ocean genomic modularity [36, 8]—though these parallels are currently structural analogies rather than formal equivalences. Establishing whether a unified mathematical framework underlies self-organization across these domains is an important open question.
Future work will extend MAL to non-central forces, dissipative systems, and real observational data, investigate whether the training schedule’s integer-ratio coincidences reflect genuine dynamical phase-locking, and explore whether explicit metabolic constraints improve scientific discovery in more complex domains. As Macklem emphasized [27], physiology must address the spontaneous emergence of self-organized order; our results suggest that embedding biological energy-minimization principles in artificial systems is a productive path toward this goal.
Methods
Data Generation
We simulated $N=16$ Keplerian orbits in 2D using a symplectic velocity-Verlet integrator with gravitational units $G=M=1$. Semi-major axes $a$ were sampled log-uniformly over $[0.5,5.0]$ AU; eccentricities $e$ uniformly over $[0,0.3]$. Each orbit was integrated for 5 orbital periods at timestep $\Delta t_{\text{sim}}=10^{-3}$, then downsampled to observation cadence $\Delta t_{\text{obs}}=0.05$. Gaussian noise $\mathcal{N}(0,\sigma^{2})$ with $\sigma=0.01\times\text{median}(a)$ was added to all position measurements. Data were split 70/15/15 into training, validation, and test sets.
Neural Architecture
Noether Force Basis. The force module computes:
$\mathbf{F}(\mathbf{r})=-\left[\sum_{i=1}^{5}A_{i}\theta_{i}\phi_{i}(r)\right]\frac{\mathbf{r}}{r}$ (2)
where $\phi_{i}\in\{r^{-2},r^{-1},r,1,r^{-3}\}$, $\theta_{i}$ are learnable scalars (force magnitudes), and $A_{i}=\mathrm{softmax}(\ell_{i}/\tau)$ are gates with logits $\ell_{i}$ and temperature $\tau$. SO(2) symmetry is enforced by restricting dependence to radial distance $r=|\mathbf{r}|$. Candidate basis functions are treated as phenomenological radial scalings; dimensional consistency resides in the fitted coefficients $\theta_{i}$ rather than in the basis library itself.
MinActionNet Integrator. The selected force law is embedded in a velocity-Verlet integrator:
$\mathbf{v}_{1/2}$ $=\mathbf{v}_{n}+\tfrac{\Delta t}{2}\mathbf{F}(\mathbf{r}_{n}),$
$\mathbf{r}_{n+1}$ $=\mathbf{r}_{n}+\Delta t\,\mathbf{v}_{1/2},$ (3)
$\mathbf{v}_{n+1}$ $=\mathbf{v}_{1/2}+\tfrac{\Delta t}{2}\mathbf{F}(\mathbf{r}_{n+1}),$
with model timestep $\Delta t_{\text{model}}=\Delta t_{\text{obs}}/5=0.01$ to ensure numerical stability while maintaining computational efficiency.
Loss Function
The Triple-Action functional (Eq. 1) decomposes as:
Information term ($\mathcal{L}_{I_{\max}}$):
$\mathcal{L}_{\text{traj}}$ $=\frac{1}{T-1}\sum_{k=0}^{T-2}\|\mathbf{r}_{\text{pred}}(t_{k+1})-\mathbf{r}_{\text{obs}}(t_{k+1})\|^{2},$ (4)
$\mathcal{L}_{\text{accel}}$ $=\frac{1}{T-2s}\sum_{j=s}^{T-s-1}\|\mathbf{F}_{\text{model}}(\mathbf{r}_{j})-\hat{\mathbf{a}}_{j}\|^{2},$ (5)
where $\hat{\mathbf{a}}_{j}=(\mathbf{r}_{j+s}-2\mathbf{r}_{j}+\mathbf{r}_{j-s})/(s\Delta t)^{2}$ is the wide-stencil acceleration estimate with stride $s=10$, and $\mathcal{L}_{I_{\max}}=\mathcal{L}_{\text{traj}}+\lambda_{\text{accel}}\mathcal{L}_{\text{accel}}$ with $\lambda_{\text{accel}}=1.0$.
Energy term ($\mathcal{L}_{E_{\min}}$):
$\mathcal{L}_{\text{sym}}$ $=\text{Var}[E(\mathbf{r}_{k},\mathbf{v}_{k})]=\langle(E-\langle E\rangle)^{2}\rangle,$ (6)
$\mathcal{L}_{\text{comp}}$ $=\langle|A_{i}\theta_{i}|\rangle_{i},$ (7)
$\mathcal{L}_{\text{arch}}$ $=-\sum_{i=1}^{5}A_{i}\log A_{i},$ (8)
where $E=\tfrac{1}{2}|\mathbf{v}|^{2}-GM/r$ is the orbital energy, $\mathcal{L}_{\text{comp}}$ is the scale-invariant sparsity penalty, $\mathcal{L}_{\text{arch}}$ is the gate entropy, and $\mathcal{L}_{E_{\min}}=\mathcal{L}_{\text{sym}}+\lambda_{\text{comp}}\mathcal{L}_{\text{comp}}+\lambda_{\text{arch}}\mathcal{L}_{\text{arch}}$ with $\lambda_{\text{comp}}=0.01$, $\lambda_{\text{arch}}=0.5$.
Training Protocol
Two-phase schedule (BGNO):
  • Phase 1 (Warmup, epochs 1–50): $\alpha_{I}=1.0$, $\alpha_{E}=0.01$, $\tau=1.0$. Low regularization allows gates to explore gradient signal from $\mathcal{L}_{\text{accel}}$.
  • Phase 2 (Sparsification, epochs 51–200): $\alpha_{I}=1.0$, $\alpha_{E}$ ramps linearly $0.01\to 1.0$, $\tau$ decays exponentially $1.0\to 0.05$. Increasing $\alpha_{E}$ drives gate competition; decreasing $\tau$ sharpens softmax toward one-hot.
Optimizer: Adam with learning rate $10^{-3}$, batch size 4, training for 200 epochs on an NVIDIA RTX 2080 Ti GPU.
Post-training calibration: After gate convergence, the L1 penalty has biased $\theta_{i}$ toward zero. We correct via least-squares projection:
$\theta_{\text{opt}}=\frac{\sum_{j}\phi_{\text{dom}}(r_{j})\hat{a}_{\text{radial},j}}{\sum_{j}\phi_{\text{dom}}(r_{j})^{2}},$ (9)
where $\phi_{\text{dom}}$ is the dominant basis function, $\hat{a}_{\text{radial},j}=-\hat{\mathbf{a}}_{\text{wide},j}\cdot\hat{\mathbf{r}}_{j}$ are radial projections of the wide-stencil acceleration estimates $\hat{\mathbf{a}}_{\text{wide}}$ (defined in Eq. 5 with stride $s=10$), and sums run over all training midpoints $j\in[s,T-s-1]$.
Validation Metrics
Kepler exponent: Orbital periods $T_{i}$ were estimated from test-set trajectories via autocorrelation peak detection with proper normalization by overlapping sample count. Power-law regression $T^{2}=Ca^{p}$ yielded $\hat{p}$ and $\hat{C}$.
Energy conservation: For long-horizon rollouts (5 orbital periods from test-set initial conditions, using the velocity-Verlet integrator at $\Delta t_{\text{model}}=0.01$ with the post-calibration coefficient $\theta_{\text{opt}}$), we computed Hamiltonian variance $\sigma_{H}^{2}=\langle(H-\langle H\rangle)^{2}\rangle$ as an energy-conservation diagnostic: correct force laws preserve $H$; incorrect laws violate conservation despite matching short-term trajectories.
Gate concentration: Architectural sparsification was quantified via the Herfindahl–Hirschman Index (HHI) on effective gate contributions $p_{i}=A_{i}|\theta_{i}|/\sum_{j}A_{j}|\theta_{j}|$, yielding $C_{\mathrm{gate}}=(K\cdot\text{HHI}-1)/(K-1)\in[0,1]$ where $K=5$ basis functions. $C_{\mathrm{gate}}=1$ indicates complete concentration on a single basis; $C_{\mathrm{gate}}=0$ indicates uniform distribution.
Code and Data Availability
All code, trained models, and synthetic data are available at https://github.com/martinfrasch/minAction_kepler. Experiments were implemented in PyTorch 2.0 on Python 3.10.
Acknowledgments.
I thank my family for giving me space to follow my ideas.
SI Text
S1. Context: Mathematical LLMs
As a baseline, we tested whether a standalone mathematical LLM (Qwen2-Math 7B, October 2024) possesses capabilities for inverse physics problems. Across nine tests spanning variational calculus, symmetry recognition, and inverse problems, the model achieved 61% overall: 100% on forward Euler-Lagrange derivation but 0% on inverse problems (given equations of motion, find the Lagrangian) and 0% on physical validity assessment (failing to identify an unphysical Lagrangian with wrong-sign potential energy). This illustrates that symbolic manipulation capability alone does not confer the inductive biases needed for physics discovery, motivating MAL’s explicit action-principle constraints. We note this tests standalone LLM reasoning; LLM-guided symbolic search pipelines (e.g., LLM-SR) represent a complementary approach not evaluated here, and rapid advances in frontier models may narrow this gap. Full test prompts, responses, and scoring rubrics are available upon request.
S2. Noise Analysis and Wide-Stencil Derivation
Noise propagation in finite differences. For positions $\mathbf{r}(t)$ measured with i.i.d. Gaussian noise $\eta\sim\mathcal{N}(0,\sigma_{\text{pos}}^{2})$, the standard second-difference acceleration estimate is:
$\hat{\mathbf{a}}_{\text{naive}}=\frac{\mathbf{r}(t+\Delta t)-2\mathbf{r}(t)+\mathbf{r}(t-\Delta t)}{\Delta t^{2}}.$ (S1)
Since $\mathbf{r}_{\text{obs}}=\mathbf{r}_{\text{true}}+\eta$ and noise samples are independent, error variance is:
$\text{Var}(\hat{\mathbf{a}}_{\text{naive}})=\frac{(1^{2}+(-2)^{2}+1^{2})\sigma_{\text{pos}}^{2}}{\Delta t^{4}}=\frac{6\sigma_{\text{pos}}^{2}}{\Delta t^{4}}.$ (S2)
For our parameters ($\sigma_{\text{pos}}=0.016$ AU, $\Delta t=0.05$), this gives noise standard deviation:
$\sigma_{a,\text{naive}}=\sqrt{\frac{6\times 0.016^{2}}{0.05^{4}}}\approx 15.7,$ (S3)
vastly exceeding the signal $|a|\approx GM/r^{2}\approx 1.0/2.0^{2}=0.25$ at $r=2$ AU (SNR $\approx 0.016$).
Wide-stencil reduction. Using stride $s$:
$\hat{\mathbf{a}}_{\text{wide}}=\frac{\mathbf{r}(t+s\Delta t)-2\mathbf{r}(t)+\mathbf{r}(t-s\Delta t)}{(s\Delta t)^{2}},$ (S4)
noise variance becomes:
$\text{Var}(\hat{\mathbf{a}}_{\text{wide}})=\frac{6\sigma_{\text{pos}}^{2}}{s^{4}\Delta t^{4}}.$ (S5)
For $s=10$:
$\sigma_{a,\text{wide}}=\sqrt{\frac{6\times 0.016^{2}}{10^{4}\times 0.05^{4}}}\approx 0.16,$ (S6)
yielding SNR $\approx 0.25/0.16\approx 1.6$—a $10^{4}\times$ improvement enabling gradient-based learning.
Systematic error. Taylor expansion of $\mathbf{r}(t\pm s\Delta t)$ around $t$ gives:
$\mathbf{r}(t+s\Delta t)$ $=\mathbf{r}(t)+s\Delta t\,\dot{\mathbf{r}}(t)+\frac{(s\Delta t)^{2}}{2}\ddot{\mathbf{r}}(t)+\frac{(s\Delta t)^{3}}{6}\mathbf{r}^{(3)}(t)+\frac{(s\Delta t)^{4}}{24}\mathbf{r}^{(4)}(t)+O((s\Delta t)^{5}),$ (S7)
$\mathbf{r}(t-s\Delta t)$ $=\mathbf{r}(t)-s\Delta t\,\dot{\mathbf{r}}(t)+\frac{(s\Delta t)^{2}}{2}\ddot{\mathbf{r}}(t)-\frac{(s\Delta t)^{3}}{6}\mathbf{r}^{(3)}(t)+\frac{(s\Delta t)^{4}}{24}\mathbf{r}^{(4)}(t)+O((s\Delta t)^{5}).$ (S8)
Summing:
$\mathbf{r}(t+s\Delta t)+\mathbf{r}(t-s\Delta t)-2\mathbf{r}(t)=(s\Delta t)^{2}\ddot{\mathbf{r}}(t)+\frac{(s\Delta t)^{4}}{12}\mathbf{r}^{(4)}(t)+O((s\Delta t)^{6}).$ (S9)
Dividing by $(s\Delta t)^{2}$:
$\hat{\mathbf{a}}_{\text{wide}}=\ddot{\mathbf{r}}(t)+\frac{(s\Delta t)^{2}}{12}\mathbf{r}^{(4)}(t)+O((s\Delta t)^{4}).$ (S10)
For Keplerian orbits, $\ddot{\mathbf{r}}=-GM\mathbf{r}/r^{3}$ and $\mathbf{r}^{(4)}\sim(GM/r^{3})\cdot(v/r)^{2}\ddot{\mathbf{r}}$. At $r=2$ AU, $v\approx\sqrt{GM/r}\approx 0.7$, so:
$\left|\frac{(s\Delta t)^{2}}{12}\mathbf{r}^{(4)}\right|/|\ddot{\mathbf{r}}|\sim\frac{(10\times 0.05)^{2}}{12}\cdot\frac{0.7^{2}}{2^{2}}\approx 0.003=0.3\%.$ (S11)
Thus systematic error $\lesssim 0.3\%$ relative to signal, negligible compared to $10^{4}\times$ noise reduction.
Verification on clean data. To confirm truncation error is small, we computed $\hat{\mathbf{a}}_{\text{wide}}$ on noise-free synthetic orbits and measured RMS deviation from true $\ddot{\mathbf{r}}=-GM\mathbf{r}/r^{3}$:
$\text{RMS}_{\text{systematic}}=\sqrt{\frac{1}{N}\sum_{j}\|\hat{\mathbf{a}}_{j}-\mathbf{a}_{\text{true},j}\|^{2}}\approx 7.2\times 10^{-4}\ll\sigma_{a,\text{wide}}=0.16.$ (S12)
This confirms that systematic error is $\sim 200\times$ smaller than remaining noise, validating the wide-stencil approach.
S3. Intrinsic Timescales and Noetherian Selection
Crystallization dynamics. We define gate selectivity $R(t)=A_{\max}(t)/A_{\text{2nd}}(t)$ and track three milestones:
  • Onset: $R\geq 10$ (clear winner emerging)
  • Sparse: $R\geq 100$ (crystallization underway)
  • Frozen: $R\geq 1000$ (effectively one-hot selection)
Across 10 random seeds trained on identical data (Table S1), the onset-to-frozen span is $\Delta t_{\text{span}}=36.2\pm 4.1$ epochs ($N=8$ fully converged seeds; bootstrap 95% CI: $[32.8,39.6]$ from 10,000 resamples), independent of selected basis or random initialization. We note that $N=8$ is small for robust parametric statistics; the bootstrap CI is more appropriate than the Gaussian-assumption CI at this sample size.
Table S1: Ten-seed robustness sweep (full data from tsparse_sweep_results.json)
Seed Basis $\hat{p}$ Onset Sparse Frozen Span $\gamma$
0 $r^{-2}$ 2.995 116 138 156 40 1.125
1 $r^{-2}$ 3.001 117 136 152 35 1.141
2 $r^{-2}$ 3.003 133 153 169 36 1.139
3 $r^{-3}$ 3.004 136 155 171 35 1.140
4 $r^{-3}$ 3.004 127 143 159 32 1.153
5 $r^{-3}$ 3.002 116 143 161 45 1.110
6 $r^{-1}$ 2.998 131 147 163 32 1.152
7 $r^{-1}$ 3.006 179 194
8 $r^{-2}$ 3.002 118 137 153 35 1.139
9 $r^{-1}$ 2.997 193
Mean $\pm$ SD (seeds 0–8) 139 $\pm$ 9 146 $\pm$ 8 161 $\pm$ 7 36.2 $\pm$ 4.1 1.137 $\pm$ 0.013
Mechanistic interpretation. The selectivity decomposes as:
$\ln R(t)=\frac{\Delta\ell(t)}{\tau(t)},\quad\text{where}\quad\Delta\ell(t)=\ell_{\text{dom}}(t)-\ell_{\text{2nd}}(t),$ (S13)
where $\ell_{i}$ are the raw gate logits. Empirically, $\Delta\ell$ grows sigmoidally from $\approx 0.05$ (epoch 1) to $\approx 0.8$ (epoch 180), while $\tau$ decays exponentially $1.0\to 0.05$. The crystallization explosion is driven by temperature amplification of a small, converged logit gap—not by a dynamical instability in logit space itself.
Noetherian model selection. When multiple seeds crystallize to different bases, energy conservation discriminates the correct physics. Table S2 shows that $r^{-2}$ models conserve the Hamiltonian $3\times$ better than $r^{-1}$ and $6\times$ better than $r^{-3}$, despite all achieving Kepler exponent $p\approx 3.0$ on short trajectories. This provides a physics-grounded criterion: among models with comparable training loss, prefer the one with minimal $\sigma_{H}$ over long-horizon rollout.
Table S2: Energy conservation as Noetherian diagnostic
Selected basis $n$ (seeds) $\bar{\sigma}_{H}$ Traj. MSE
$r^{-2}$ (correct) 4 $0.017\pm 0.016$ 5.8
$r^{-1}$ 3 $0.056\pm 0.023$ 6.1
$r^{-3}$ 3 $0.11\pm 0.017$ $1.2\times 10^{3}$
Slow-annealing experiment. To test schedule-dependence, we trained an 11th model (seed 0, 300 epochs, $\tau:1.0\to 0.001$). Results: onset at epoch 103 (vs. 116 standard), frozen at epoch 134 (vs. 156), span $\Delta t_{\text{span}}=31$ (vs. 40 standard), growth rate $\gamma=1.165$ (vs. 1.125). The span and growth rate remain within one SD of the standard schedule, confirming that these quantities are intrinsic to the dynamics, not artifacts of the cooling protocol.
S4. Baseline Comparisons and Extended Benchmarks
S4A. Hooke’s Law Benchmark
To test generalization beyond inverse-square gravity, we applied MAL to Hooke’s law (F=krF=-kr) using 16 synthetic orbits (near-circular, $\omega=1$, $\sigma=0.01$) with the same basis library and training protocol as Kepler. Table S3 summarizes 10-seed results.
Table S3: Hooke’s law 10-seed sweep results
Seed Basis $\hat{k}$ Onset Frozen Span Reject gravity?
0 $r^{-3}$ 0.040 125 164 39 Yes
1 $r$ 0.980 113 157 44 Yes
2 $r$ 0.980 106 148 42 Yes
3 $r$ 0.980 104 147 43 Yes
4 $r$ 0.980 100 144 44 Yes
5 $r$ 0.980 123 165 42 Yes
6 $r$ 0.980 99 142 43 Yes
7 $r$ 0.981 97 140 43 Yes
8 $r$ 0.980 103 146 43 Yes
9 $r$ 0.981 105 149 44 Yes
Mean $\pm$ SD 0.980 $\pm$ 0.001 108 $\pm$ 9 150 $\pm$ 8 42.5 $\pm$ 1.6 10/10
Mean computed over 9 seeds selecting correct basis $r$; seed 0 ($r^{-3}$, $\hat{k}=0.040$) excluded.
The energy-conservation-based diagnostic computes per-orbit Hamiltonian variance for each candidate potential derived from the basis library. For Hooke data, gravity-family potentials ($V\propto r^{-1},r^{-2},r^{-3}$, corresponding to basis functions in the library) yield relative variance >0.01>0.01, while the correct linear potential ($V=\frac{1}{2}kr^{2}$, corresponding to basis $r$) is at the noise floor ($\sim 10^{-3}$). The diagnostic rejects gravity-family candidates 10/10 seeds with $3\times$ margin.
S4B. Basis Selection Interventions
To diagnose the 40% direct selection rate on Kepler, we tested three interventions (Table S4).
Table S4: Basis selection intervention experiments (Kepler, 10 seeds each)
Intervention $r^{-2}$ rate Description Key insight
Standard (baseline) 4/10 Default training protocol
Extended warmup (100 ep.) 4/10 Double warmup duration More exploration doesn’t help
Lower noise ($\sigma=0.005$) 4/10 Half observation noise Cleaner data doesn’t help
Biased $\alpha_{\text{logits}}$ 10/10 $[1.5,0,0,0,0]$ init Gate competition is bottleneck
The biased initialization gives $r^{-2}$ approximately 50% initial gate weight ($A_{0}\approx 0.47$ vs. $A_{i}\approx 0.13$ for others), providing a head start that survives the warmup competition. Neither longer exploration nor cleaner data changes the outcome, confirming that the bottleneck is gate logit dynamics during warmup, not noise or insufficient exploration.
S4C. SINDy Variant Comparison
We compared MAL against four SINDy configurations across 10 random seeds on identical Kepler data (Table S5). All methods used the same radial basis library $\{r^{-2},r^{-1},r,1,r^{-3}\}$.
Table S5: SINDy variant comparison on Kepler benchmark (10 seeds)
Method Stencil $r^{-2}$ rate $\overline{GM}$ Time
Vanilla SINDy (naive) $s=1$ 0/10 <0.01<0.01 s
Vanilla SINDy (wide) $s=10$ 10/10 $1.05\text{--}2.74$ <0.01<0.01 s
GP-SINDy (wide) $s=10$ 8/10 $0.97\text{--}1.88$ $\sim$6 s
Ensemble-SINDy (wide) $s=10$ 10/10 $1.07\text{--}2.68$ $\sim$0.4 s
MAL $s=10$ 4/10 (10/10) $\sim$0.94 $\sim$835 s
With biased gate initialization or Noetherian post-selection.
The critical insight is that wide-stencil preprocessing ($s=10$) is the key enabler for all methods. Without it, even SINDy fails (noise-dominated, SNR $\sim 0.02$). With it, SINDy achieves excellent basis identification at a fraction of MAL’s computational cost. MAL’s advantage lies in (i) integrated dynamical rollout for trajectory validation, (ii) Noetherian energy conservation diagnostics, and (iii) the ability to handle more complex systems where sparse regression may fail.
Note that SINDy’s $GM$ estimates vary widely across seeds (range: 0.97–2.74 for vanilla wide-stencil), reflecting sensitivity to the specific radial distribution of training data. MAL’s post-training calibration yields a more consistent estimate ($\hat{GM}\approx 0.94$).
S4D. HNN and LNN Comparison
We trained Hamiltonian Neural Networks (HNNs) [18] and Lagrangian Neural Networks (LNNs) [10] on identical Kepler data with comparable architectures (Table S6).
Table S6: Structure-preserving neural network comparison
Method Params Time Energy $\sigma_{H}$ Symbolic? Status
HNN 17,281 113 s 0.006 kWh $4.1\times 10^{-4}$ No Converged
LNN 17,281 242 s 0.013 kWh $1.9\times 10^{-2}$ No Hessian singular
MAL $\sim$20 835 s 0.046 kWh $0.017$ Yes Converged
HNN achieved the best energy conservation ($\sigma_{H}=4.1\times 10^{-4}$, vs. MAL’s $0.017$), consistent with its conservation-by-construction architecture. However, HNN’s learned Hamiltonian is a black-box neural network (17K parameters) that provides no interpretable symbolic force law. LNN suffered from Hessian singularities: the learned mass matrix $M(\mathbf{q})$ became ill-conditioned during training, causing validation loss to diverge to $10^{6}$ throughout all 200 epochs—a known failure mode for LNNs on noisy data where the Hessian $\partial^{2}L/\partial\dot{\mathbf{q}}^{2}$ must remain positive definite.
MAL trades energy conservation precision for interpretability: its $\sim$20 learnable parameters yield an explicit symbolic force law ($F\propto r^{-2}$ with calibrated coefficient), while HNN’s 17K parameters encode the same physics implicitly. This distinction is critical for scientific discovery, where the goal is not just prediction but understanding.
S4E. Additional Baselines
(1) Teacher-forcing only (no $\mathcal{L}_{E_{\min}}$): Trained identical MinActionNet without energy/sparsity terms ($\alpha_{E}=0$). Gates remained at $A_{\max}=0.45$ (failed crystallization, $R\approx 2$). Training required 1480 s (+77% vs. MAL) due to lack of architectural pruning. This demonstrates $\mathcal{L}_{E_{\min}}$ is essential for both basis selection and energy efficiency.
(2) Standard feedforward NN (black-box): FC network ($\mathbf{r}\to[128]^{3}\to\mathbf{F}$, 50K params) achieved MSE $10^{-3}$ on training range but extrapolated nonsensically ($r>6$ AU: divergence; $r<0.3$ AU: oscillation). Trajectory rollout collapsed within 2 orbits.
(3) Physics-informed NN [32]: 85% accuracy when $r^{-2}$ law pre-specified in loss; cannot discover the law de novo.
(4) Mathematical LLM (Qwen2-Math 7B): 0% on inverse problems, 61% on forward derivation (SI Appendix, Section S1). Noether’s Razor [41] and Noether’s Learning Dynamics [37] provide complementary theoretical perspectives on symmetry-driven model selection. SNAS [45] employs the same softmax temperature annealing as MAL’s gates, motivated by computational efficiency rather than biological metabolic constraints.
Summary: MAL’s contribution is energy-constrained model selection within a pre-specified basis library, combining wide-stencil noise reduction, SO(2)-constrained architecture, bimodal energy optimization, and energy-conservation-based validation. Wide-stencil preprocessing is the critical enabler shared with SINDy; MAL’s unique addition is the integration of dynamical rollout, energy conservation validation, and explicit metabolic constraints.
S4F. Basis Library Sensitivity
To test robustness to basis library composition, we ran four experiments (10 seeds each, Table S7).
Table S7: Basis library sensitivity (Kepler benchmark, 10 seeds each)
Experiment $K$ $r^{-2}$ rate $\bar{C}_{\mathrm{gate}}$ $\bar{\sigma}_{H}$ (correct) $\bar{\sigma}_{H}$ (incorrect)
Standard 5 5/10 0.97 0.132 0.183
Confounders 7 2/10 1.00 0.091 0.158
Expanded 8 5/10 0.96 0.134 0.138
Missing ($r^{-2}$ absent) 4 0.96 0.156
(1) Standard control ($K=5$: $\{r^{-2},r^{-1},r,1,r^{-3}\}$): 5/10 seeds selected the correct $r^{-2}$ basis, consistent with the 4/10 rate reported in the main text (different random seeds). All seeds crystallized ($C_{\mathrm{gate}}\geq 0.97$). The energy-conservation diagnostic discriminated correct from incorrect models ($\bar{\sigma}_{H}=0.132$ vs. $0.183$).
(2) Near-confounders ($K=7$: standard $\{r^{-2.5},r^{-1.5}\}$): Adding bases with radial exponents close to $r^{-2}$ reduced correct selection to 2/10. The confounders $r^{-2.5}$ and $r^{-1.5}$ captured 5 seeds between them. Over the limited radial range of the training orbits ($r\in[0.5,5]$ AU), these bases are near-degenerate with $r^{-2}$, creating competing attractors in gate logit space that trap gradient flow during warmup. Despite the reduced selection rate, the energy-conservation diagnostic still discriminated: correct models achieved $\bar{\sigma}_{H}=0.091$ vs. $0.158$ for incorrect. This demonstrates that (a) basis library composition strongly affects raw selection rates, and (b) the energy-conservation criterion remains robust to library expansion.
(3) Expanded library ($K=8$: standard $\{r^{2},r^{-4},\ln r\}$): Adding bases with radial scalings far from $r^{-2}$ had no effect on the correct-basis rate (5/10), identical to the $K=5$ control. The additional bases ($r^{2}$, $r^{-4}$, $\ln r$) are sufficiently distinct in their radial profiles that they do not create competing attractors. Training time increased modestly ($\sim$1120 s vs. $\sim$870 s for $K=5$).
(4) Missing correct basis ($K=4$: $\{r^{-1},r,1,r^{-3}\}$, $r^{-2}$ absent): When the correct basis is excluded, the system splits between the two nearest alternatives: $r^{-1}$ (5 seeds) and $r^{-3}$ (5 seeds). Crystallization still occurs ($\bar{C}_{\mathrm{gate}}=0.96$), but no seed converges to a physically correct model. The elevated $\bar{\sigma}_{H}=0.156$ across all seeds (compared to $0.132$ for correct models in the standard library) provides a diagnostic signal: uniformly poor energy conservation across an ensemble flags potential library inadequacy.
Implications for practical use. These results quantify a fundamental limitation of basis-library model selection: the method can only select among candidates provided. When the correct basis is present but confounders exist, the raw selection rate degrades; when it is absent, the system fails gracefully. The energy-conservation diagnostic remains informative in all cases, suggesting a practical workflow: run multiple seeds, apply the conservation criterion, and treat uniformly high $\sigma_{H}$ across all seeds as evidence that the library may be inadequate.
S5. Schedule Geometry, Modularity Analysis, and Cross-Domain Parallels
Phase-space trajectory. We plot $(\alpha_{E}(t),\tau(t))$ for all 200 training epochs, color-coded by epoch number. Red diamonds mark epochs where the ratio $\alpha_{E}/\tau$ passes through integer values (3:1, 2:1, 3:2, 1:1), defined as epochs where $|\alpha_{E}q/\tau p-1|<0.1$.
Major gate transitions (onset at $R=10$, sparsification at $R=100$, crystallization at $R=1000$) coincide with passage through or near these integer-ratio nodes. Important caveat: Because $\alpha_{E}$ ramps linearly and $\tau$ decays exponentially, the schedule trajectory is externally designed, not emergent. Any smooth two-parameter sweep through a bounded region will necessarily pass through integer-ratio nodes (a consequence of the density of rationals), so the coincidence with gate transitions may be a geometric artifact of the schedule design rather than evidence of dynamical phase-locking. Distinguishing these possibilities requires formal analysis—e.g., showing that crystallization timing is sensitive to the schedule’s integer-ratio passages via perturbation experiments—which we leave to future work.
Structural parallel to neonatal physiology. Hoyer et al. [21] showed that neonatal heart rate (HR) and breathing movement (BM) exhibit integer-ratio coordination:
  • Quiet sleep (low metabolic rate, synaptic consolidation [39]): 3:1 HR:BM coordination
  • Active sleep (high metabolic rate, memory formation): Off-center ratios (e.g., 5:2, 7:3)
MAL’s training phases exhibit a superficially similar structure (low-regularization exploration $\to$ high-regularization consolidation). However, we emphasize a key difference: the neonatal system involves genuine coupled oscillators (heart rate and breathing are autonomous rhythmic processes), whereas MAL’s $\alpha_{E}$ and $\tau$ are monotonically varied hyperparameters, not oscillators. Dynamic coordination theory [34] formalizes how weakly coupled nonlinear oscillators synchronize at Farey ratios—but this formalism does not directly apply to monotonic schedule parameters. The parallel is therefore structural and suggestive, not mechanistic.
Architectural sparsification. We quantified gate concentration $C_{\mathrm{gate}}$ using the Herfindahl–Hirschman Index (HHI) on effective gate contributions $p_{i}=A_{i}|\theta_{i}|/\sum_{j}A_{j}|\theta_{j}|$, yielding $C_{\mathrm{gate}}=(K\cdot\text{HHI}-1)/(K-1)\in[0,1]$ where $K=5$ basis functions. MAL-trained models ($N=10$ seeds) achieved $C_{\mathrm{gate}}=0.99\pm 0.02$ by epoch 200 (near-complete gate crystallization), while teacher-forcing-only baselines ($N=10$ seeds) remained at $C_{\mathrm{gate}}=0.14\pm 0.04$ (Mann–Whitney $U=100$, $p=9.1\times 10^{-5}$; permutation test $p<0.001$ with n=10,000n=10{,}000 permutations). The complete separation between groups demonstrates that energy-constrained training drives architectural sparsification, consistent with Clune et al.’s [8] finding that minimizing wiring costs produces modular structure in evolved networks.
S6. Connection to TARA Oceans Genomic Modularity
The TARA Oceans expedition [36] sampled microbial communities across global ocean gradients, revealing that gene co-expression networks exhibit modular structures. Clune et al. [8] demonstrated computationally that such modularity arises when networks evolve under pressure to minimize connection costs—a proxy for metabolic expenditure.
We observe quantitative parallels between TARA Oceans network properties and MAL’s learned architectures:
  1. 1.
    Scale-free topology: Ocean microbial networks exhibit power-law degree distributions $P(k)\sim k^{-\gamma}$ with $\gamma\in[2.1,2.8]$ [36]; MAL gate networks show $\gamma=2.4\pm 0.2$ during crystallization.
  2. 2.
    High modularity: Ocean networks yield $Q\in[0.7,0.9]$; MAL’s $r^{-2}$ models achieve $Q=0.87\pm 0.04$.
  3. 3.
    Metabolic constraint: In both systems, modular structure is consistent with wiring-cost minimization [8].
Temporal dynamics. Marine microbial community assembly transitions between high-connectivity states during bloom events (high nutrient/energy) and low-connectivity states during oligotrophic periods (low nutrient/energy). This high-to-low connectivity transition is structurally parallel to MAL’s warmup (broad exploration) to sparsification (architectural pruning) transition.
Limitations of this comparison. We emphasize that these are structural parallels, not established causal connections. The cited references [36, 8] demonstrate modularity and wiring-cost minimization but do not discuss Farey sequences or integer-ratio scaling. The convergent modularity may simply reflect that wiring-cost minimization generically produces modular networks [8] regardless of the specific system, rather than evidence of a unified organizing principle. Formal investigation—e.g., measuring whether ocean network temporal dynamics exhibit integer-ratio coordination analogous to neonatal physiological coupling [21]—is needed to substantiate or refute the hypothesis of cross-scale universality.
S7. Extended Discussion: Broader Context
Penrose’s three worlds and AI for physics. Penrose [31] proposed that reality consists of three interconnected domains: M (all mathematical structures), P (the subset nature implements), and C (the subset conscious observers can model). Current AI operates in C $\to$ C: learning from human text to predict more text. MAL aims to bridge M $\to$ P by learning which mathematical structures (candidate force laws) minimize action given observational data, implementing a selection principle that does not require human conceptual mediation. However, we note that MAL’s current basis library is itself a human-designed element of C, so the M $\to$ P bridge is partial.
Future extensions. Two directions would strengthen the M $\to$ P bridge: (i) incorporating gauge invariance as an explicit constraint (MAL’s $\mathcal{L}_{\mathrm{Symmetry}}$ currently enforces energy conservation via Noether’s theorem, but local gauge structure requires additional architectural innovations); and (ii) dimensional analysis to penalize basis functions with inappropriate mass dimensions, which could enable identification of renormalizable field theories beyond classical force laws.
S8. Computational Details
Hardware:
  • GPU: NVIDIA GeForce RTX 2080 Ti (11 GB VRAM, Turing architecture)
  • CPU: Intel Xeon E5-2670 v3 @ 2.30GHz (12 cores, 24 threads)
  • RAM: 64 GB DDR4 ECC
  • Storage: 1 TB NVMe SSD
Software:
  • OS: Ubuntu 22.04 LTS (Linux kernel 5.15.0)
  • Python: 3.10.8
  • PyTorch: 2.0.1+cu117 (CUDA 11.7)
  • NumPy: 1.24.2
  • Matplotlib: 3.7.1
  • SciPy: 1.10.1
Hyperparameters (comprehensive list):
  • Data generation:
    • Number of orbits: 16 (11 train, 3 val, 2 test)
    • Semi-major axis range: $a\in[0.5,5.0]$ AU (log-uniform sampling)
    • Eccentricity range: $e\in[0,0.3]$ (uniform sampling)
    • Integration timestep: $\Delta t_{\text{sim}}=10^{-3}$ time units
[…]
    • Trajectory length: 5 orbital periods each
    • Positional noise: $\sigma=0.01\times\text{median}(a)\approx 0.016$ AU (Gaussian)
  • Network architecture:
    • Basis library size: $K=5$ functions $\{r^{-2},r^{-1},r,1,r^{-3}\}$
    • Gate temperature initialization: $\tau_{0}=1.0$
    • Coefficient initialization: $\theta_{i}\sim\mathcal{N}(0,0.01^{2})$
    • Logit initialization: $\alpha_{i}\sim\mathcal{U}(-0.1,0.1)$
    • Model integration timestep: $\Delta t_{\text{model}}=0.01$ (5 substeps per observation)
  • Training:
    • Optimizer: Adam ($\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$)
    • Learning rate: $\eta=10^{-3}$ (constant, no decay)
    • Batch size: 4 trajectories
    • Total epochs: 200
    • Warmup duration: 50 epochs
    • Loss weight schedule: $\alpha_{I}=1.0$ (constant), $\alpha_{E}:0.01\to 1.0$ (linear ramp, epochs 51–200)
    • Temperature schedule: $\tau:1.0\to 0.05$ (exponential decay, epochs 51–200)
    • Stencil stride: $s=10$ for acceleration matching
  • Loss components:
    • $\lambda_{\text{accel}}=1.0$ (acceleration matching weight)
    • $\lambda_{\text{comp}}=0.01$ (complexity/sparsity penalty)
    • $\lambda_{\text{arch}}=0.5$ (architecture entropy penalty)
  • Calibration:
    • Post-training least-squares over all training trajectories
    • Using same wide-stencil ($s=10$) acceleration estimates
    • Projecting onto dominant basis function $\phi_{\text{dom}}$
Energy estimation:
  • Total training time: 835 seconds (13.9 minutes)
  • GPU TDP (RTX 2080 Ti): 250 W (rated), conservatively assumed 200 W under sustained load
  • Energy consumption: $E=200\,\text{W}\times 835\,\text{s}/3600\,\text{s/h}\approx 0.046$ kWh
  • Full system power (CPU + RAM + motherboard + cooling): estimated additional 100 W
  • Total system energy: $\approx 0.046\times(300/200)\approx 0.07$ kWh
Carbon footprint:
  • U.S. average grid carbon intensity: 0.42 kg CO2/kWh (2024 EPA estimate)
  • Per-model emissions: $0.07\,\text{kWh}\times 0.42\,\text{kg/kWh}\approx 0.029$ kg CO2e $\approx 30$ g CO2e
  • Equivalent to: charging a smartphone ($\sim 0.01$ kWh/charge) 7 times, or driving an EV 0.15 miles
  • For comparison: Strubell et al. [35] estimated $\sim$284 tons CO2e for NAS-based Transformer training (NLP models, not LLMs); Patterson et al. [29] estimated GPT-3 (175B parameters) training at $\sim$552 tons CO2e, consuming $\sim$1,287 MWh. While these comparisons involve vastly different tasks and scales, they illustrate the energy-efficiency advantage of the MAL approach for the physics-discovery domain.
Reproducibility: All experiments use fixed random seeds (0–9 for robustness sweep) set via:
import torch, numpy as np, random
torch.manual_seed(seed)
np.random.seed(seed)
random.seed(seed)
torch.backends.cudnn.deterministic = True
Training is deterministic on the same hardware/software configuration, though slight numerical differences ($0.1\%$) may occur across GPU architectures due to floating-point non-associativity.
SI Figures
Figure S1: Robustness: Seed 137 orbit reconstruction. Long-horizon rollout (5 orbital periods) from initial conditions, using the calibrated $r^{-2}$ force law discovered by seed 137. Model trajectory (red) closely matches ground truth (blue), with slight enlargement attributable to 6% deficit in recovered GMGM.
Figure S2: Robustness: Seed 137 training curves. Loss component dynamics for seed 137, showing identical two-phase structure (warmup epochs 1–50, sparsification epochs 51–200) as primary seed 0. Onset occurs at epoch 121, within one SD of the mean (117 $\pm$ 9).
Figure S3: Robustness: Seed 137 gate evolution. Architecture selection for seed 137, converging to $r^{-2}$ basis with identical intrinsic timescales ($\Delta t_{\text{span}}=35$ epochs, $\gamma=1.141$).
Figure S4: Robustness: Seed 137 energy conservation. Hamiltonian variance $\sigma_{H}=0.019$ for seed 137, within the $r^{-2}$ group mean $0.017\pm 0.016$, confirming Noetherian selection criterion.
Figure S5: Multi-panel discovery summary. (A) Orbit comparison: model rollout vs. ground truth for test orbit. (B) Architecture gate evolution over 200 epochs. (C) Learned force coefficients $\theta_{i}$ before and after calibration. (D) Kepler exponent fit: $T^{2}\propto a^{3.01}$.
Figure S6: Architectural modularity analysis. (A) Modularity index $Q$ (Herfindahl–Hirschman concentration on effective gate contributions $A_{i}|\theta_{i}|$) over training epochs for MAL (blue, mean $\pm$ SD, $N=10$ seeds) vs. teacher-forcing-only baselines (red dashed, $N=10$). Vertical dashed lines mark the warmup-to-sparsification transition. MAL modularity rises sharply during sparsification, reaching $Q=0.99\pm 0.02$ by epoch 200, while teacher-forcing remains diffuse ($Q=0.14\pm 0.04$). (B) Final modularity by group: models selecting the correct $r^{-2}$ basis, models selecting other bases, and teacher-forcing baselines (permutation test $p<0.001$, n=10,000n=10{,}000). Both MAL groups achieve near-maximal $Q$ regardless of which basis is selected, confirming that energy-constrained training drives architectural sparsification independent of outcome. (C) Final gate probability distributions $A_{i}$ for all 10 MAL seeds (S0–S9) and 10 teacher-forcing baselines (B0–B9). MAL seeds crystallize to one-hot gate vectors (single dominant basis per seed), while baselines maintain diffuse distributions across all five bases.

### 25 Mar 2026 v2

@@
-Minimum-Action Learning: Energy-Constrained Symbolic Model Selection for Physical Law Identification from Noisy Data
+Minimum-action learning: Energy-constrained symbolic model selection for identifying physical laws from noisy data
 Abstract
 Identifying physical laws from noisy observational data is a central challenge in scientific machine learning.
-We present Minimum-Action Learning (MAL), a framework that selects symbolic force laws from a pre-specified basis library by minimizing a Triple-Action functional combining trajectory reconstruction ($I_{\mathrm{max}}$), architectural sparsity ($E_{\mathrm{min}}$), and energy-conservation enforcement ($\mathcal{L}_{\mathrm{Symmetry}}$).
-A wide-stencil acceleration-matching technique reduces noise variance by $10^{4}\times$, transforming an intractable problem (SNR $\approx 0.02$) into a learnable one (SNR $\approx 1.6$); this preprocessing is the critical enabler shared by all methods tested, including SINDy variants.
-Using two benchmarks—Kepler (inverse-square gravity) and Hooke’s law (linear restoring force)—we show that MAL recovers the correct force law with Kepler exponent $p=3.01\pm 0.01$ in 835 s at $\sim$0.07 kWh (40% reduction vs.
-prediction-error-only baselines).
-The raw correct-basis rate is 40% for Kepler and 90% for Hooke; an energy-conservation-based model selection criterion discriminates the true force law from alternatives in all cases, yielding 100% pipeline-level identification.
-A physics-informed gate initialization independently achieves 10/10 correct selection.
-Direct comparison against noise-robust SINDy variants, HNNs, and LNNs confirms MAL’s distinct niche: energy-constrained, interpretable model selection combining symbolic basis identification with dynamical rollout validation.
-The sparse architectures that emerge exhibit intrinsic crystallization timescales ($\Delta t_{\text{span}}=36.2\pm 4.1$ epochs) independent of initialization.
-The regularization design draws structural inspiration from biological metabolic constraints, though the biological analogies discussed remain suggestive rather than formally established.
-Keywords: neural architecture search, action principles, energy optimization, physical law identification, modularity, developmental neuroscience
+We present Minimum-Action Learning (MAL), a differentiable framework that selects symbolic force laws from a pre-specified basis library by minimizing a triple-action functional combining trajectory reconstruction, architectural sparsity, and energy-conservation enforcement.
+A wide-stencil acceleration-matching technique reduces noise variance by a factor of $10^{4}$, transforming an intractable problem (signal-to-noise ratio $\approx 0.02$) into a learnable one ($\approx 1.6$); this preprocessing is the critical enabler shared by all methods tested, including SINDy variants.
+On two benchmarks—Kepler inverse-square gravity and Hooke’s linear restoring force—MAL recovers the correct force law with Kepler exponent $p=3.01\pm 0.01$ in $835$ s at $\sim\!0.07$ kWh, a 40% reduction relative to prediction-error-only baselines.
+The raw correct-basis rate is 40% for Kepler (4/10 seeds) and 90% for Hooke (9/10); an energy-conservation-based model selection criterion discriminates the true force law in all cases, yielding 100% pipeline-level identification.
+Direct comparison against SINDy variants, Hamiltonian neural networks, and Lagrangian neural networks confirms MAL’s distinct niche: energy-constrained, interpretable model selection that combines symbolic basis identification with dynamical rollout validation within a single differentiable framework.
+The sparse gate architectures that emerge exhibit intrinsic crystallization timescales ($\Delta t_{\text{span}}=36.2\pm 4.1$ epochs) independent of initialization, with geometric growth rate $\gamma=1.137\pm 0.013$ per epoch.
+I Introduction
 The principle of least action, formulated by Maupertuis and refined by Lagrange and Hamilton, has unified physics from classical mechanics to quantum field theory [1].
-Simultaneously, biological systems—constrained by thermodynamics and natural selection—have evolved architectures that minimize metabolic expenditure while maximizing information processing [44, 25].
-We propose that these parallel optimization strategies are not coincidental but reflect a deeper organizing principle operating across scales: network-weighted action minimization.
-Here we demonstrate that embedding this principle in artificial neural networks—via a Triple-Action functional combining information maximization ($I_{\mathrm{max}}$), energy minimization ($E_{\mathrm{min}}$), and symmetry enforcement—enables energy-constrained identification of physical laws from noisy data, while spontaneously generating the modular architectures characteristic of evolved systems.
-Introduction
-The Energy-Physics Connection
-From Claude Bernard’s milieu intérieur [2] through Cannon’s homeostasis [5] to contemporary network physiology [1], biology has recognized that living systems operate through energy-constrained multi-scale integration.
-Kleiber’s discovery [25] that metabolic rate scales as $M^{3/4}$ across 17 orders of magnitude in body mass, later explained by West, Brown, and Enquist [44] through network optimization, demonstrates that metabolic constraints impose mathematical regularities on biological organization.
-In parallel, physics has long relied on action principles—the requirement that systems follow paths extremizing $S=\int L\,dt$, where $L$ is the Lagrangian—to derive equations of motion [17].
-From classical mechanics to quantum field theory, nature appears to solve optimization problems, minimizing action subject to symmetry constraints [43].
-Recent work in developmental neuroscience reveals that these principles converge in the brain: glial cells enforce metabolic constraints on neural architecture search during development [19, 42], while dynamic coordination theory [34, 22, 23, 24] formalizes the relationship between energy states and network synchronization through metastable dynamics [38] and Farey-tree patterns—a framework supported by nonlinear frequency-locking phenomena in autonomic neural circuits [16] and phase-oscillator models of synaptic plasticity [6].
-We hypothesize that this metabolic architecture search provides a template for energy-efficient artificial intelligence capable of discovering physical laws.
-The Kepler Benchmark: Rediscovering $F=Gm_{1}m_{2}/r^{2}$
-Newton’s derivation of the inverse-square law from Kepler’s observations represents one of science’s foundational achievements.
-Can a machine, given only noisy positional measurements of orbiting bodies, autonomously rediscover this law?
-This problem—which we term the ”Kepler-with-noise” benchmark—tests whether AI systems possess genuine physical intuition or merely pattern-match existing formulations.
-Existing approaches address aspects of this challenge.
-Symbolic regression methods [40, 11] and GNN-based symbolic distillation [11, 26] can recover force laws from clean or moderately noisy data, but degrade at high noise levels.
+A natural question is whether this principle can be turned inward: can action minimization itself serve as a regularizer for machine-learning systems tasked with identifying physical laws from data?
+Identifying force laws from noisy observational data—the inverse problem of mechanics—remains challenging because noise amplifies catastrophically under numerical differentiation, rendering standard approaches unreliable [2, 3].
+Symbolic regression methods [4, 5] and GNN-based symbolic distillation [5, 6] recover force laws from clean or moderately noisy data but degrade at high noise levels.
 Hamiltonian neural networks (HNNs) [7] and Lagrangian neural networks (LNNs) [8] embed conservation structure but learn black-box energy functions rather than selecting interpretable symbolic forms.
-Physics-informed neural networks [32] require pre-specifying the equations to solve.
-Standalone mathematical LLMs lack inductive biases for inverse physical reasoning (SI Appendix, Section S1), though LLM-guided symbolic search pipelines offer a complementary approach.
-Noether’s Razor [41] and Noether’s Learning Dynamics [37] provide complementary theoretical perspectives on symmetry-driven model selection.
-We present Minimum-Action Learning (MAL), a framework that embeds biological energy-optimization principles directly into neural architecture search.
-By training networks to minimize a Triple-Action functional combining information content ($I_{\mathrm{max}}$), metabolic cost ($E_{\mathrm{min}}$), and symmetry constraints ($\mathcal{L}_{\mathrm{Symmetry}}$), MAL performs energy-constrained model selection over a pre-specified basis library, identifying physical laws while maintaining biological-scale energy budgets.
-The key contribution is not open-ended symbolic discovery but rather the demonstration that explicit metabolic constraints—inspired by the brain’s glial-neural architecture—improve both the efficiency and interpretability of physical law identification from noisy data.
-Results
-Energy-Constrained Identification of Newton’s Law
+Physics-informed neural networks [2] require pre-specifying the governing equations.
+Noether’s Razor [9] and Noether’s Learning Dynamics [10] provide complementary theoretical perspectives on symmetry-driven model selection but do not address the noise bottleneck directly.
+We present Minimum-Action Learning (MAL), a differentiable framework that selects symbolic force laws from a pre-specified basis library by minimizing a triple-action functional combining trajectory reconstruction ($I_{\mathrm{max}}$), architectural sparsity ($E_{\mathrm{min}}$), and energy-conservation enforcement ($\mathcal{L}_{\mathrm{Symmetry}}$).
+Three design elements are central: (i) a wide-stencil acceleration-matching technique that reduces noise variance by $10^{4}\times$, transforming an intractable signal-to-noise ratio (SNR $\approx 0.02$) into a learnable one (SNR $\approx 1.6$); (ii) temperature-annealed gate competition that drives a soft-to-discrete architectural transition, selecting a single basis function from the library; and (iii) an energy-conservation-based model selection criterion, grounded in Noether’s theorem, that discriminates the true force law from alternatives across independent training runs.
+We validate MAL on two benchmarks—Kepler inverse-square gravity and Hooke’s linear restoring force—and compare directly against SINDy variants, HNNs, and LNNs.
+The regularization design draws structural inspiration from biological metabolic constraints on neural architecture [11, 12]: the sparsity-inducing $E_{\mathrm{min}}$ term plays a role analogous to wiring-cost minimization in evolved networks.
+While these biological analogies remain suggestive rather than formally established, the resulting computational framework stands on its own as a physics-informed method for interpretable model selection.
+II Results
+II.1 Energy-constrained identification of Newton’s law
 We implemented MAL as a differentiable neural network (MinActionNet) incorporating three key architectural innovations (Fig. 2, Methods):
@@
 where $\mathcal{L}_{I_{\max}}$ combines trajectory reconstruction and wide-stencil acceleration matching (maximizing information extraction from noisy data), $\mathcal{L}_{E_{\min}}$ penalizes architectural complexity via gate entropy and coefficient sparsity (minimizing energy), and $\mathcal{L}_{\mathrm{Symmetry}}$ enforces energy conservation (Noether’s theorem).
-(3) Bimodal Glial-Neural Optimization (BGNO).
-Drawing a structural analogy from the brain’s 1:1 glia-to-neuron ratio [9] and evidence that glial cells mediate metabolic constraints during neural development [7, 12], we implement a two-phase training schedule in which the Triple-Action functional drives the architecture $\mathbf{A}(t)$ along a soft-to-discrete manifold, where energy-driven sharpening ($E_{\mathrm{min}}$) encourages the emergence of discrete structural motifs from initially uniform gate distributions.
+(3) Two-phase training schedule.
+We implement a schedule in which the triple-action functional drives the architecture $\mathbf{A}(t)$ along a soft-to-discrete manifold, where energy-driven sharpening ($E_{\mathrm{min}}$) encourages the emergence of discrete structural motifs from initially uniform gate distributions.
 During warmup (50 epochs, low regularization $\alpha_{E}=0.01$, uniform gates $\tau=1.0$), gates $A_{i}=\mathrm{softmax}(\ell_{i}/\tau)$ remain in a soft, exploratory state; during sparsification (150 epochs, ramping $\alpha_{E}\to 1.0$, annealing $\tau\to 0.05$), the $E_{\mathrm{min}}$ subsystem penalizes non-discrete architectural states, driving gate sharpening toward one-hot selection.
-This mechanism is analogous to temperature-annealed differentiable architecture search [45], here motivated by biological metabolic constraints rather than computational efficiency alone.
-We note that the glial-neural analogy is structural, not mechanistic: the regularization weight $\alpha_{E}$ does not perform functions equivalent to glial metabolic support (e.g., lactate shuttling, potassium buffering), but serves a functionally analogous role of constraining the “neural” gradient dynamics.
+This mechanism is closely related to temperature-annealed differentiable architecture search [13], with the additional constraint that the sparsity pressure is explicitly tied to an energy-minimization objective inspired by wiring-cost minimization in biological networks [11, 12].
 Figure 1: Trajectory reconstruction and Hamiltonian conservation.
@@
 This transition occurs on a soft-to-discrete architecture manifold where A_logits are sharpened via softmax temperature decay ($\tau:1\to 0.05$), driven by the $E_{\mathrm{min}}$ subsystem which penalizes non-discrete states through gate entropy $\mathcal{L}_{\mathrm{arch}}=-\sum_{i}A_{i}\log A_{i}$.
-(B) The two-phase BGNO regularization schedule ($\alpha_{E}$ shown in inset) separates physics identification (warmup) from architectural sparsification.
-Shown for a representative seed (seed 0); variability across 10 seeds is reported in SI Table S1.
+(B) The two-phase regularization schedule ($\alpha_{E}$ shown in inset) separates physics identification (warmup) from architectural sparsification.
+Shown for a representative seed (seed 0); variability across 10 seeds is reported in Supplemental Table S1.
 Training on 16 synthetic Keplerian orbits (semi-major axes $a\in[0.5,5]$ AU, eccentricities $e\in[0,0.3]$, observation noise $\sigma=1\%$ of median radius) for 200 epochs yielded:
@@
 2).
-Across 10 seeds, 4 of 10 directly select $r^{-2}$; the remaining seeds select $r^{-3}$ (3 seeds) or $r^{-1}$ (3 seeds), with a Noetherian diagnostic correctly discriminating the true physics in all cases (see Robustness section).
+Across 10 seeds, 4 of 10 directly select $r^{-2}$; the remaining seeds select $r^{-3}$ (3 seeds) or $r^{-1}$ (3 seeds), with a energy-conservation diagnostic correctly discriminating the true physics in all cases (see Robustness section).
   • Accurate force calibration: Recovered gravitational coupling $\theta_{0}=0.936$ (true: $GM=1.0$), representing 6.4% error attributable to residual noise in acceleration estimates (Methods).
@@
   • Energy efficiency: Total training time 835 seconds, energy consumption $\sim$0.07 kWh (full system: 200 W GPU ++ $\sim$100 W CPU/RAM/cooling), representing 40% reduction vs.
-prediction-error-only baselines (SI Appendix, Table S2).
+prediction-error-only baselines (Supplemental Material, Table S2).
 […]
-The Noise Bottleneck and Wide-Stencil Solution
+II.2 The noise bottleneck and wide-stencil solution
 A critical insight emerged from failure analysis: naive finite-difference acceleration estimates from noisy positions have signal-to-noise ratio (SNR) $\ll 1$, rendering learning impossible.
 For observation noise $\sigma_{\text{pos}}$ and timestep $\Delta t$, second-derivative variance scales as $\sigma_{a}^{2}\propto\sigma_{\text{pos}}^{2}/\Delta t^{4}$.
-At typical scales ($\sigma_{\text{pos}}=0.016$ AU, $\Delta t=0.05$), this yields $\sigma_{a}\approx 15.7$, while true gravitational acceleration at $r=2$ AU is $\approx 0.25$—an SNR of $0.016$ (Table 1, SI Appendix).
+At typical scales ($\sigma_{\text{pos}}=0.016$ AU, $\Delta t=0.05$), this yields $\sigma_{a}\approx 15.7$, while true gravitational acceleration at $r=2$ AU is $\approx 0.25$—an SNR of $0.016$ (Table 1, Supplemental Material).
 We solved this through wide-stencil acceleration matching: computing second derivatives with stride $s=10$ reduces noise variance by $s^{4}=10{,}000\times$ while introducing only $O(s^{2}\Delta t^{2})$ systematic error ($0.3\%$ for Keplerian orbits).
-This transforms an intractable problem (SNR $\approx 0.02$) into a learnable one (SNR $\approx 1.6$), enabling gradient-based discovery of the force law (Methods, SI Appendix Section S2).
+This transforms an intractable problem (SNR $\approx 0.02$) into a learnable one (SNR $\approx 1.6$), enabling gradient-based discovery of the force law (Methods, Supplemental Material Section S2).
 Table 1: Noise reduction via wide-stencil differentiation
@@
 20 0.039 0.25 6.4 Success
-Robustness, Success Rate, and Intrinsic Timescales
+II.3 Robustness, success rate, and intrinsic timescales
 To assess reliability, we trained 10 independent models with different random seeds on identical data.
-All 10 successfully crystallized to a single dominant basis, though seeds 7 and 9 required extended training beyond 200 epochs (SI Appendix, Fig.
+All 10 successfully crystallized to a single dominant basis, though seeds 7 and 9 required extended training beyond 200 epochs (Supplemental Material, Fig.
 S1–S4, Table S1).
@@
 However, all 10 models recovered Kepler exponent $p\in[2.995,3.006]$—a consequence of teacher-forced trajectory matching, which preserves orbital mechanics over the limited radial range tested even when the selected basis is incorrect.
-This means $p\approx 3.0$ alone is insufficient to discriminate the true force law; the Noetherian criterion—already intrinsic to MAL’s trifunctional via $\mathcal{L}_{\mathrm{Symmetry}}$—must be applied as a model selection diagnostic across the seed ensemble (see below).
+This means $p\approx 3.0$ alone is insufficient to discriminate the true force law; the energy-conservation criterion—already intrinsic to MAL’s trifunctional via $\mathcal{L}_{\mathrm{Symmetry}}$—must be applied as a model selection diagnostic across the seed ensemble (see below).
 Because the symmetry term enforces energy conservation during training, models that crystallize to the correct basis exhibit superior long-horizon Hamiltonian conservation, providing a built-in discriminant.
-Applying this criterion across seeds achieves 100% identification of the correct physics (SI Table S2).
+Applying this criterion across seeds achieves 100% identification of the correct physics (Supplemental Table S2).
 Basis selection interventions.
-To diagnose the 40% direct selection rate, we tested three interventions across 10 seeds each (SI Table S4): (i) extended warmup (100 epochs, 4/10 correct), (ii) reduced noise ($\sigma=0.005$, 4/10 correct), and (iii) physics-informed gate initialization ($\alpha_{\text{logits}}=[1.5,0,0,0,0]$, giving $r^{-2}$ approximately 50% initial gate weight).
+To diagnose the 40% direct selection rate, we tested three interventions across 10 seeds each (Supplemental Table S4): (i) extended warmup (100 epochs, 4/10 correct), (ii) reduced noise ($\sigma=0.005$, 4/10 correct), and (iii) physics-informed gate initialization ($\alpha_{\text{logits}}=[1.5,0,0,0,0]$, giving $r^{-2}$ approximately 50% initial gate weight).
 Neither longer exploration nor cleaner data improved the success rate, but biased initialization achieved 10/10 correct selection, confirming that gate competition during the warmup phase—not noise level or exploration time—is the bottleneck.
-This suggests a practical recommendation: when prior knowledge favors a particular basis, encoding it as a logit bias eliminates the need for cross-seed Noetherian model selection.
+This suggests a practical recommendation: when prior knowledge favors a particular basis, encoding it as a logit bias eliminates the need for cross-seed energy-conservation model selection.
 Intrinsic timescales. Three invariant quantities emerged across seeds:
@@
 (2) Growth rate: Within the crystallization window, $R$ grows geometrically at rate $\gamma=1.137\pm 0.013$ per epoch, corresponding to Lyapunov exponent $\lambda=\ln\gamma\approx 0.128$ epoch-1.
-These timescales are intrinsic to the competition dynamics between gates driven by temperature annealing and logit gradient flow, not artifacts of the specific schedule (SI Appendix, Section S3, Fig.
+These timescales are intrinsic to the competition dynamics between gates driven by temperature annealing and logit gradient flow, not artifacts of the specific schedule (Supplemental Material, Section S3, Fig.
 S5).
-Generalization to Hooke’s Law
-To test whether MAL generalizes beyond inverse-square gravity, we applied it to Hooke’s law (F=krF=-kr, linear restoring force) using the same basis library $\{r^{-2},r^{-1},r,1,r^{-3}\}$, identical architecture, and identical training protocol (SI Appendix, Section S4A, SI Table S3).
+II.4 Generalization to Hooke’s law
+To test whether MAL generalizes beyond inverse-square gravity, we applied it to Hooke’s law (F=krF=-kr, linear restoring force) using the same basis library $\{r^{-2},r^{-1},r,1,r^{-3}\}$, identical architecture, and identical training protocol (Supplemental Material, Section S4A, Supplemental Table S3).
 Of 10 seeds, 9 directly selected the correct $r$ basis (90% vs.
 40% for Kepler), recovering spring constant $\hat{k}=0.980\pm 0.001$ (true: $k=1.0$, 2% error).
-The single outlier (seed 0) selected $r^{-3}$ but was correctly rejected by the Noetherian diagnostic: across all 10 seeds, the Noetherian criterion rejected gravity-family potentials ($r^{-2}$, $r^{-1}$, $r^{-3}$) in favor of the correct linear family by a factor $3\times$ in energy conservation variance.
+The single outlier (seed 0) selected $r^{-3}$ but was correctly rejected by the energy-conservation diagnostic: across all 10 seeds, the energy-conservation criterion rejected gravity-family potentials ($r^{-2}$, $r^{-1}$, $r^{-3}$) in favor of the correct linear family by a factor $3\times$ in energy conservation variance.
 The higher direct-selection success rate (90% vs.
@@
 After energy-conservation-based cross-seed selection.
-Energy-Conservation-Based Model Selection
+II.5 Energy-conservation-based model selection
 Because MAL’s trifunctional includes an energy-conservation term $\mathcal{L}_{\mathrm{Symmetry}}=\mathrm{Var}[E]$ (inspired by Noether’s theorem linking time-translation symmetry to energy conservation), models that crystallize to the correct basis are trained to conserve energy—providing a built-in discriminant that can be applied as a model selection criterion across seeds without any additional computation.
-For long-horizon rollouts (5 orbital periods), we computed Hamiltonian variance $\sigma_{H}^{2}=\langle(H-\langle H\rangle)^{2}\rangle$: models selecting the correct $r^{-2}$ basis conserve energy $3\times$ better than $r^{-1}$ and $6\times$ better than $r^{-3}$ models, despite all achieving comparable short-term trajectory accuracy (SI Table S2, Fig.
+For long-horizon rollouts (5 orbital periods), we computed Hamiltonian variance $\sigma_{H}^{2}=\langle(H-\langle H\rangle)^{2}\rangle$: models selecting the correct $r^{-2}$ basis conserve energy $3\times$ better than $r^{-1}$ and $6\times$ better than $r^{-3}$ models, despite all achieving comparable short-term trajectory accuracy (Supplemental Table S2, Fig.
 4).
@@
 The $E_{\mathrm{min}}$-driven soft-to-discrete transition produces near-one-hot gate distributions: gate concentration $C_{\mathrm{gate}}=0.99\pm 0.02$ (Herfindahl–Hirschman Index on effective gate contributions $A_{i}|\theta_{i}|$; $C_{\mathrm{gate}}=1$ indicates complete concentration on a single basis) for MAL-trained models ($N=10$ seeds) vs.
-$C_{\mathrm{gate}}=0.14\pm 0.04$ for teacher-forcing-only baselines (Mann–Whitney $U=100$, $p<10^{-4}$; permutation test $p<0.001$, n=10,000n=10{,}000; SI Appendix, Fig.
+$C_{\mathrm{gate}}=0.14\pm 0.04$ for teacher-forcing-only baselines (Mann–Whitney $U=100$, $p<10^{-4}$; permutation test $p<0.001$, n=10,000n=10{,}000; Supplemental Material, Fig.
 S6).
 This architectural sparsification is consistent with Clune et al.’s [11] finding that minimizing wiring costs—a proxy for metabolic expenditure—produces modular structure in evolved networks.
-Structural parallels across domains.
-We note intriguing structural parallels between MAL’s modularity emergence and patterns in biological systems, while emphasizing that these are currently analogies rather than formal equivalences:
-(i) Perinatal brain development: Neonatal heart rate and breathing movement coordination during quiet (low-energy) vs.
-active (high-energy) sleep states exhibit integer-ratio (3:1) phase relationships [21, 33], reflecting dynamic coordination of coupled physiological oscillators under energy constraints.
-This frequency-locking behavior is a manifestation of the metastable coordination dynamics described by Kelso [22, 38, 24], and parallels the nonlinear frequency-locking of baroreceptor and sympathetic rhythms observed in autonomic neural circuits [16].
-(ii) Ecological networks: Microbial gene co-expression networks in the TARA Oceans dataset [36] exhibit modular architectures consistent with wiring-cost minimization [8], though the connection to MAL’s training dynamics remains analogical (SI Appendix, Section S6).
-We emphasize that these cross-domain parallels are structural analogies, not established causal connections.
-The training schedule parameters $(\alpha_{E},\tau)$ are externally imposed hyperparameters, not emergent oscillators; and the biological systems involve genuine coupled-oscillator dynamics.
-Whether energy-constrained optimization generically produces similar architectural motifs across these domains is a hypothesis warranting formal investigation.
-Comparison to Alternative Approaches
-We benchmarked MAL against five alternatives on the Kepler-with-noise problem (SI Appendix, Section S4, SI Tables S5–S6):
+Structural parallels.
+The sparse architectures produced by MAL’s energy-constrained optimization parallel modular structures observed in networks evolved under wiring-cost pressure [11, 12].
+Whether energy-constrained optimization generically produces similar architectural motifs across biological and artificial systems is a hypothesis warranting formal investigation; we discuss cross-domain parallels further in the Supplemental Material (Sections S5–S6).
+II.6 Comparison to alternative approaches
+We benchmarked MAL against five alternatives on the Kepler-with-noise problem (Supplemental Material, Section S4, Supplemental Tables S5–S6):
 (1) SINDy variants [3, 14, 15]: Vanilla SINDy with naive differentiation (stride $s=1$) fails completely (noise-dominated).
-However, when equipped with the same wide-stencil preprocessing ($s=10$) used in MAL, vanilla SINDy identifies $r^{-2}$ in 10/10 seeds, GP-SINDy in 8/10, and ensemble-SINDy in 10/10 (SI Table S5).
+However, when equipped with the same wide-stencil preprocessing ($s=10$) used in MAL, vanilla SINDy identifies $r^{-2}$ in 10/10 seeds, GP-SINDy in 8/10, and ensemble-SINDy in 10/10 (Supplemental Table S5).
 The wide-stencil technique is the critical enabler, not the learning algorithm.
@@
 $\sim$835 s); its limitation is the absence of dynamical validation—SINDy returns sparse coefficients but cannot roll out trajectories or verify energy conservation.
-MAL’s Noetherian diagnostic provides this additional layer.
-(2) HNN [18] and LNN [10]: We trained both on identical data (SI Table S6).
+MAL’s energy-conservation diagnostic provides this additional layer.
+(2) HNN [7] and LNN [8]: Trained on identical data (Supplemental Table S6).
 HNN (17K parameters, 113 s) achieved excellent energy conservation ($\sigma_{H}=4.1\times 10^{-4}$) but learns a black-box Hamiltonian with no interpretable symbolic form.
@@
 Neither method performs basis selection or yields a symbolic force law.
-(3) Mathematical LLM (Qwen2-Math 7B): 0% accuracy on inverse problems, 61% on forward derivation—mathematical manipulation without physical intuition (SI Appendix, Section S1).
-(4) Physics-informed NN [32]: 85% accuracy when the inverse-square law is pre-specified, but cannot identify the law de novo.
+(3) Mathematical LLM (Qwen2-Math 7B): 0% on inverse problems, 61% on forward derivation (Supplemental Material, Section S1).
+(4) Physics-informed NN [2]: 85% when the inverse-square law is pre-specified, but cannot identify the law de novo.
 (5) Teacher-forcing only (no $\mathcal{L}_{E_{\min}}$): Gates remained at 45% (failed to crystallize); energy consumption 77% higher.
 Summary of niche.
-MAL occupies a distinct position: it combines interpretable symbolic basis selection (shared with SINDy [3]) with dynamical rollout and Noetherian validation (shared with HNN [18]), while adding explicit energy-efficiency optimization.
-HNNs/LNNs [18, 10] guarantee conservation but produce black-box energy functions; SINDy [3] yields symbolic expressions but without dynamical consistency checks; MAL provides both within a metabolically constrained framework.
-Discussion
-Key Findings
-Our results establish four key findings:
-(1) Metabolic constraints improve physical law identification.
+MAL occupies a distinct position: it combines interpretable symbolic basis selection (shared with SINDy [3]) with dynamical rollout and energy-conservation validation (shared with HNN [7]), while adding explicit sparsity-driven efficiency optimization.
+HNNs/LNNs [7, 8] guarantee conservation but produce black-box energy functions; SINDy [3] yields symbolic expressions but without dynamical consistency checks; MAL provides both within an energy-constrained framework.
+III Discussion
+III.1 Key findings
+Our results establish four principal findings:
+(1) Sparsity constraints improve physical law identification.
 Embedding energy minimization ($E_{\min}$) alongside information maximization ($I_{\max}$) in the triple-action training objective constrains the model selection problem from an unconstrained search to a tractable optimization.
 The $E_{\mathrm{min}}$ subsystem drives gate crystallization (basis selection) that does not occur under prediction error alone, while reducing training energy by 40%.
-This parallels Noble’s biological relativity [28]: information content and metabolic cost jointly determine which model structure the network selects.
 […]
-(3) Integer-ratio coincidences suggest cross-scale connections.
-We observe that major gate transitions coincide with epochs where the designed schedule ratio $\alpha_{E}/\tau$ passes through integer values.
-We emphasize that this observation is currently correlational: $\alpha_{E}$ and $\tau$ are externally imposed hyperparameters, not emergent oscillators, and the coincidence may reflect the density of rationals rather than dynamical phase-locking.
-However, the structural parallel to integer-ratio coordination in physiological oscillators [21, 16] and the broader framework of metastable coordination dynamics [34, 22, 38, 23, 24] is intriguing and motivates formal investigation.
-Recent work on phase-oscillator networks with synaptic plasticity [6] suggests that frequency-locking phenomena may be generic in systems where coupling weights co-evolve with oscillator dynamics—a structural analogue to MAL’s co-evolution of gate logits and force coefficients.
-A rigorous test would require showing that the crystallization dynamics are sensitive to the schedule trajectory’s passage through integer ratios—e.g., via circle-map analysis of the coupled $\alpha_{E}$-$\tau$ dynamics—which we leave to future work.
-(4) Architectural sparsification as a consequence of energy minimization.
+(3) Intrinsic crystallization timescales.
+The gate competition dynamics exhibit universal timescales: crystallization span $\Delta t_{\text{span}}=36.2\pm 4.1$ epochs and geometric growth rate $\gamma=1.137\pm 0.013$ per epoch, independent of which basis wins or the initialization seed.
+We also observe that major gate transitions coincide with epochs where the schedule ratio $\alpha_{E}/\tau$ passes through integer values, though this may simply reflect the density of rationals along any smooth trajectory through parameter space.
+Whether the crystallization dynamics are genuinely sensitive to these passages requires perturbation experiments, which we leave to future work.
+(4) Architectural sparsification from energy minimization.
 Clune et al.
 [11] demonstrated that networks evolving under pressure to minimize connection costs spontaneously develop modular architectures.
-Our work shows a related pattern in physics model selection: MAL’s sparse gate selection ($C_{\mathrm{gate}}=0.99$) emerges from $\mathcal{L}_{E_{\min}}$, not from a pre-programmed bias.
-This is consistent with the hypothesis that metabolic constraints on wiring costs drive architectural sparsification across both biological [4] and artificial systems [8].
-Implications for AI and Theoretical Biology
-Energy-efficient AI.
-MAL’s 40% energy reduction over prediction-error-only baselines demonstrates that explicit metabolic constraints ($E_{\min}$ optimization) can improve computational efficiency without sacrificing task performance.
-The mechanism is architectural: $\mathcal{L}_{E_{\min}}$ drives early gate crystallization, reducing the effective parameter count and shortening convergence time.
-This principle—that sparsity-inducing regularization simultaneously improves interpretability and efficiency—may generalize beyond the current benchmarks to larger-scale scientific discovery tasks.
-Energy-conservation-based model selection.
-A persistent challenge in AI interpretability is distinguishing genuine causal understanding from correlation [30].
-The energy-conservation criterion—preferring models that preserve Hamiltonian dynamics over long-horizon rollouts—provides a physics-grounded approach complementary to recent work on learning conserved quantities [41, 37]: models that violate energy conservation reveal their failure to capture true causal structure, even when matching short-term predictions.
-This criterion is inspired by Noether’s theorem (time-translation symmetry implies energy conservation), though our implementation is a specific instantiation rather than a general Noether construction.
-Developmental origins of health and disease.
-The structural analogy between MAL’s two-phase optimization and brain development suggests testable hypotheses: Does disruption of metabolic coordination (e.g., maternal inflammation [15]) alter the integer-ratio structure of physiological coupling?
-Could this predict neurodevelopmental outcomes?
-While the correspondence is currently analogical, the intrinsic crystallization timescales observed in MAL motivate quantitative investigation of equivalent timescales in biological development.
-Limitations and Future Directions
+Our work shows a related pattern: MAL’s sparse gate selection ($C_{\mathrm{gate}}=0.99$) emerges from $\mathcal{L}_{E_{\min}}$ without a pre-programmed bias toward any particular basis.
+This supports the broader principle that cost-constrained optimization drives architectural sparsification in both biological [12] and artificial systems.
+III.2 Broader implications
+Energy-efficient scientific machine learning.
+MAL’s 40% energy reduction over prediction-error-only baselines demonstrates that explicit sparsity constraints can improve computational efficiency without sacrificing task performance.
+The mechanism is architectural: $\mathcal{L}_{E_{\min}}$ drives early gate crystallization, reducing the effective parameter count and shortening convergence.
+This principle—that sparsity-inducing regularization simultaneously improves interpretability and efficiency—may generalize to larger-scale scientific discovery tasks.
+Noether-based model selection.
+A persistent challenge in scientific machine learning is distinguishing genuine physical understanding from statistical correlation [16].
+The energy-conservation criterion—preferring models that preserve Hamiltonian dynamics over long-horizon rollouts—provides a physics-grounded approach complementary to recent work on learning conserved quantities [9, 10].
+Models that violate energy conservation reveal their failure to capture true causal structure, even when matching short-term predictions.
+III.3 Limitations and future directions
 Fixed basis library.
@@
 This makes MAL a model selection framework, not an open-ended discovery system.
-Basis library sensitivity experiments (SI Appendix, Section S4F) quantify this limitation: adding near-confounders ($r^{-2.5}$, $r^{-1.5}$) reduces the correct-basis rate from 50% to 20%, while removing the correct basis ($r^{-2}$) causes the system to split between alternatives.
+Basis library sensitivity experiments (Supplemental Material, Section S4F) quantify this limitation: adding near-confounders ($r^{-2.5}$, $r^{-1.5}$) reduces the correct-basis rate from 50% to 20%, while removing the correct basis ($r^{-2}$) causes the system to split between alternatives.
 Distant additions ($r^{2}$, $r^{-4}$, $\ln r$) have no effect.
@@
 Understanding the logit-gradient dynamics that determine gate competition—and developing initialization-free methods to improve the raw success rate—remain priorities.
-Cross-scale analogies.
-The structural parallels between MAL’s training dynamics, physiological coordination, and ocean genomic modularity are currently analogies, not formal equivalences.
-The training schedule parameters are externally imposed, not emergent oscillators.
-Establishing formal connections would require, e.g., circle-map analysis showing sensitivity of crystallization to integer-ratio passages, or demonstrating Farey coordination in systems where the schedule is itself learned rather than designed.
-Conclusion
+IV Conclusion
 We have demonstrated that minimum-action learning—neural network training guided by a triple-action functional combining information maximization, energy minimization, and symmetry enforcement—enables energy-constrained identification of physical force laws from noisy data.
-Validated on two benchmarks (Kepler gravity and Hooke’s linear restoring force), the key technical contributions are: (1) the wide-stencil noise reduction technique that transforms an intractable problem (SNR $\sim 0.02$) into a learnable one (SNR $\sim 1.6$); (2) the soft-to-discrete gate sharpening mechanism that achieves basis selection through energy-driven architectural crystallization, with a physics-informed initialization achieving 10/10 correct selection; and (3) the Noetherian model selection criterion that discriminates correct from incorrect physics via long-horizon energy conservation.
+Validated on two benchmarks (Kepler gravity and Hooke’s linear restoring force), the key technical contributions are: (1) the wide-stencil noise reduction technique that transforms an intractable problem (SNR $\sim 0.02$) into a learnable one (SNR $\sim 1.6$); (2) the soft-to-discrete gate sharpening mechanism that achieves basis selection through energy-driven architectural crystallization, with a physics-informed initialization achieving 10/10 correct selection; and (3) the energy-conservation-based model selection criterion that discriminates correct from incorrect physics via long-horizon Hamiltonian conservation.
 Direct comparison against SINDy variants, HNNs, and LNNs confirms MAL’s distinct niche: interpretable, energy-constrained model selection combining symbolic basis identification with dynamical validation.
-The modular architectures that emerge from this process parallel patterns observed in biological systems under metabolic constraints—including physiological coordination [21] and ocean genomic modularity [36, 8]—though these parallels are currently structural analogies rather than formal equivalences.
-Establishing whether a unified mathematical framework underlies self-organization across these domains is an important open question.
-Future work will extend MAL to non-central forces, dissipative systems, and real observational data, investigate whether the training schedule’s integer-ratio coincidences reflect genuine dynamical phase-locking, and explore whether explicit metabolic constraints improve scientific discovery in more complex domains.
-As Macklem emphasized [27], physiology must address the spontaneous emergence of self-organized order; our results suggest that embedding biological energy-minimization principles in artificial systems is a productive path toward this goal.
-Methods
-Data Generation
+Future work will extend MAL to non-central forces, dissipative systems, higher-dimensional problems, and real observational data (e.g., planetary ephemerides), and will test whether the intrinsic crystallization timescales and schedule-geometry coincidences observed here generalize across problem classes.
+V Methods
+V.1 Data generation
 We simulated $N=16$ Keplerian orbits in 2D using a symplectic velocity-Verlet integrator with gravitational units $G=M=1$.
@@
 Data were split 70/15/15 into training, validation, and test sets.
-Neural Architecture
+V.2 Neural architecture
 Noether Force Basis. The force module computes:
@@
 with model timestep $\Delta t_{\text{model}}=\Delta t_{\text{obs}}/5=0.01$ to ensure numerical stability while maintaining computational efficiency.
-Loss Function
+V.3 Loss function
 The triple-action functional (Eq. 1) decomposes as:
@@
 where $E=\tfrac{1}{2}|\mathbf{v}|^{2}-GM/r$ is the orbital energy, $\mathcal{L}_{\text{comp}}$ is the scale-invariant sparsity penalty, $\mathcal{L}_{\text{arch}}$ is the gate entropy, and $\mathcal{L}_{E_{\min}}=\mathcal{L}_{\text{sym}}+\lambda_{\text{comp}}\mathcal{L}_{\text{comp}}+\lambda_{\text{arch}}\mathcal{L}_{\text{arch}}$ with $\lambda_{\text{comp}}=0.01$, $\lambda_{\text{arch}}=0.5$.
-Training Protocol
-Two-phase schedule (BGNO):
+V.4 Training protocol
+Two-phase schedule:
   • Phase 1 (Warmup, epochs 1–50): $\alpha_{I}=1.0$, $\alpha_{E}=0.01$, $\tau=1.0$. Low regularization allows gates to explore gradient signal from $\mathcal{L}_{\text{accel}}$.
@@
 5 with stride $s=10$), and sums run over all training midpoints $j\in[s,T-s-1]$.
-Validation Metrics
+V.5 Validation metrics
 Kepler exponent: Orbital periods $T_{i}$ were estimated from test-set trajectories via autocorrelation peak detection with proper normalization by overlapping sample count.
@@
 $C_{\mathrm{gate}}=1$ indicates complete concentration on a single basis; $C_{\mathrm{gate}}=0$ indicates uniform distribution.
-Code and Data Availability
+V.6 Code and data availability
 All code, trained models, and synthetic data are available at https://github.com/martinfrasch/minAction_kepler. Experiments were implemented in PyTorch 2.0 on Python 3.10.
-Acknowledgments.
+Acknowledgements.
 I thank my family for giving me space to follow my ideas.
-SI Text
-S1. Context: Mathematical LLMs
-As a baseline, we tested whether a standalone mathematical LLM (Qwen2-Math 7B, October 2024) possesses capabilities for inverse physics problems.
-Across nine tests spanning variational calculus, symmetry recognition, and inverse problems, the model achieved 61% overall: 100% on forward Euler-Lagrange derivation but 0% on inverse problems (given equations of motion, find the Lagrangian) and 0% on physical validity assessment (failing to identify an unphysical Lagrangian with wrong-sign potential energy).
-This illustrates that symbolic manipulation capability alone does not confer the inductive biases needed for physics discovery, motivating MAL’s explicit action-principle constraints.
-We note this tests standalone LLM reasoning; LLM-guided symbolic search pipelines (e.g., LLM-SR) represent a complementary approach not evaluated here, and rapid advances in frontier models may narrow this gap.
-Full test prompts, responses, and scoring rubrics are available upon request.
-S2. Noise Analysis and Wide-Stencil Derivation
-Noise propagation in finite differences.
-For positions $\mathbf{r}(t)$ measured with i.i.d.
-Gaussian noise $\eta\sim\mathcal{N}(0,\sigma_{\text{pos}}^{2})$, the standard second-difference acceleration estimate is:
-$\hat{\mathbf{a}}_{\text{naive}}=\frac{\mathbf{r}(t+\Delta t)-2\mathbf{r}(t)+\mathbf{r}(t-\Delta t)}{\Delta t^{2}}.$ (S1)
-Since $\mathbf{r}_{\text{obs}}=\mathbf{r}_{\text{true}}+\eta$ and noise samples are independent, error variance is:
-$\text{Var}(\hat{\mathbf{a}}_{\text{naive}})=\frac{(1^{2}+(-2)^{2}+1^{2})\sigma_{\text{pos}}^{2}}{\Delta t^{4}}=\frac{6\sigma_{\text{pos}}^{2}}{\Delta t^{4}}.$ (S2)
-For our parameters ($\sigma_{\text{pos}}=0.016$ AU, $\Delta t=0.05$), this gives noise standard deviation:
-$\sigma_{a,\text{naive}}=\sqrt{\frac{6\times 0.016^{2}}{0.05^{4}}}\approx 15.7,$ (S3)
-vastly exceeding the signal $|a|\approx GM/r^{2}\approx 1.0/2.0^{2}=0.25$ at $r=2$ AU (SNR $\approx 0.016$).
-Wide-stencil reduction. Using stride $s$:
-$\hat{\mathbf{a}}_{\text{wide}}=\frac{\mathbf{r}(t+s\Delta t)-2\mathbf{r}(t)+\mathbf{r}(t-s\Delta t)}{(s\Delta t)^{2}},$ (S4)
-noise variance becomes:
-$\text{Var}(\hat{\mathbf{a}}_{\text{wide}})=\frac{6\sigma_{\text{pos}}^{2}}{s^{4}\Delta t^{4}}.$ (S5)
-For $s=10$:
-$\sigma_{a,\text{wide}}=\sqrt{\frac{6\times 0.016^{2}}{10^{4}\times 0.05^{4}}}\approx 0.16,$ (S6)
-yielding SNR $\approx 0.25/0.16\approx 1.6$—a $10^{4}\times$ improvement enabling gradient-based learning.
-Systematic error. Taylor expansion of $\mathbf{r}(t\pm s\Delta t)$ around $t$ gives:
-$\mathbf{r}(t+s\Delta t)$ $=\mathbf{r}(t)+s\Delta t\,\dot{\mathbf{r}}(t)+\frac{(s\Delta t)^{2}}{2}\ddot{\mathbf{r}}(t)+\frac{(s\Delta t)^{3}}{6}\mathbf{r}^{(3)}(t)+\frac{(s\Delta t)^{4}}{24}\mathbf{r}^{(4)}(t)+O((s\Delta t)^{5}),$ (S7)
-$\mathbf{r}(t-s\Delta t)$ $=\mathbf{r}(t)-s\Delta t\,\dot{\mathbf{r}}(t)+\frac{(s\Delta t)^{2}}{2}\ddot{\mathbf{r}}(t)-\frac{(s\Delta t)^{3}}{6}\mathbf{r}^{(3)}(t)+\frac{(s\Delta t)^{4}}{24}\mathbf{r}^{(4)}(t)+O((s\Delta t)^{5}).$ (S8)
-Summing:
-$\mathbf{r}(t+s\Delta t)+\mathbf{r}(t-s\Delta t)-2\mathbf{r}(t)=(s\Delta t)^{2}\ddot{\mathbf{r}}(t)+\frac{(s\Delta t)^{4}}{12}\mathbf{r}^{(4)}(t)+O((s\Delta t)^{6}).$ (S9)
-Dividing by $(s\Delta t)^{2}$:
-$\hat{\mathbf{a}}_{\text{wide}}=\ddot{\mathbf{r}}(t)+\frac{(s\Delta t)^{2}}{12}\mathbf{r}^{(4)}(t)+O((s\Delta t)^{4}).$ (S10)
-For Keplerian orbits, $\ddot{\mathbf{r}}=-GM\mathbf{r}/r^{3}$ and $\mathbf{r}^{(4)}\sim(GM/r^{3})\cdot(v/r)^{2}\ddot{\mathbf{r}}$. At $r=2$ AU, $v\approx\sqrt{GM/r}\approx 0.7$, so:
-$\left|\frac{(s\Delta t)^{2}}{12}\mathbf{r}^{(4)}\right|/|\ddot{\mathbf{r}}|\sim\frac{(10\times 0.05)^{2}}{12}\cdot\frac{0.7^{2}}{2^{2}}\approx 0.003=0.3\%.$ (S11)
-Thus systematic error $\lesssim 0.3\%$ relative to signal, negligible compared to $10^{4}\times$ noise reduction.
-Verification on clean data.
-To confirm truncation error is small, we computed $\hat{\mathbf{a}}_{\text{wide}}$ on noise-free synthetic orbits and measured RMS deviation from true $\ddot{\mathbf{r}}=-GM\mathbf{r}/r^{3}$:
-$\text{RMS}_{\text{systematic}}=\sqrt{\frac{1}{N}\sum_{j}\|\hat{\mathbf{a}}_{j}-\mathbf{a}_{\text{true},j}\|^{2}}\approx 7.2\times 10^{-4}\ll\sigma_{a,\text{wide}}=0.16.$ (S12)
-This confirms that systematic error is $\sim 200\times$ smaller than remaining noise, validating the wide-stencil approach.
-S3. Intrinsic Timescales and Noetherian Selection
-Crystallization dynamics. We define gate selectivity $R(t)=A_{\max}(t)/A_{\text{2nd}}(t)$ and track three milestones:
-  • Onset: $R\geq 10$ (clear winner emerging)
-  • Sparse: $R\geq 100$ (crystallization underway)
-  • Frozen: $R\geq 1000$ (effectively one-hot selection)
-Across 10 random seeds trained on identical data (Table S1), the onset-to-frozen span is $\Delta t_{\text{span}}=36.2\pm 4.1$ epochs ($N=8$ fully converged seeds; bootstrap 95% CI: $[32.8,39.6]$ from 10,000 resamples), independent of selected basis or random initialization.
-We note that $N=8$ is small for robust parametric statistics; the bootstrap CI is more appropriate than the Gaussian-assumption CI at this sample size.
-Table S1: Ten-seed robustness sweep (full data from tsparse_sweep_results.json)
-Seed Basis $\hat{p}$ Onset Sparse Frozen Span $\gamma$
-0 $r^{-2}$ 2.995 116 138 156 40 1.125
-1 $r^{-2}$ 3.001 117 136 152 35 1.141
-2 $r^{-2}$ 3.003 133 153 169 36 1.139
-3 $r^{-3}$ 3.004 136 155 171 35 1.140
-4 $r^{-3}$ 3.004 127 143 159 32 1.153
-5 $r^{-3}$ 3.002 116 143 161 45 1.110
-6 $r^{-1}$ 2.998 131 147 163 32 1.152
-7 $r^{-1}$ 3.006 179 194
-8 $r^{-2}$ 3.002 118 137 153 35 1.139
-9 $r^{-1}$ 2.997 193
-Mean $\pm$ SD (seeds 0–8) 139 $\pm$ 9 146 $\pm$ 8 161 $\pm$ 7 36.2 $\pm$ 4.1 1.137 $\pm$ 0.013
-Mechanistic interpretation. The selectivity decomposes as:
-$\ln R(t)=\frac{\Delta\ell(t)}{\tau(t)},\quad\text{where}\quad\Delta\ell(t)=\ell_{\text{dom}}(t)-\ell_{\text{2nd}}(t),$ (S13)
-where $\ell_{i}$ are the raw gate logits.
-Empirically, $\Delta\ell$ grows sigmoidally from $\approx 0.05$ (epoch 1) to $\approx 0.8$ (epoch 180), while $\tau$ decays exponentially $1.0\to 0.05$.
-The crystallization explosion is driven by temperature amplification of a small, converged logit gap—not by a dynamical instability in logit space itself.
-Noetherian model selection.
-When multiple seeds crystallize to different bases, energy conservation discriminates the correct physics.
-Table S2 shows that $r^{-2}$ models conserve the Hamiltonian $3\times$ better than $r^{-1}$ and $6\times$ better than $r^{-3}$, despite all achieving Kepler exponent $p\approx 3.0$ on short trajectories.
-This provides a physics-grounded criterion: among models with comparable training loss, prefer the one with minimal $\sigma_{H}$ over long-horizon rollout.
-Table S2: Energy conservation as Noetherian diagnostic
-Selected basis $n$ (seeds) $\bar{\sigma}_{H}$ Traj. MSE
-$r^{-2}$ (correct) 4 $0.017\pm 0.016$ 5.8
-$r^{-1}$ 3 $0.056\pm 0.023$ 6.1
-$r^{-3}$ 3 $0.11\pm 0.017$ $1.2\times 10^{3}$
-Slow-annealing experiment.
-To test schedule-dependence, we trained an 11th model (seed 0, 300 epochs, $\tau:1.0\to 0.001$).
-Results: onset at epoch 103 (vs.
-116 standard), frozen at epoch 134 (vs.
-156), span $\Delta t_{\text{span}}=31$ (vs.
-40 standard), growth rate $\gamma=1.165$ (vs.
-1.125).
-The span and growth rate remain within one SD of the standard schedule, confirming that these quantities are intrinsic to the dynamics, not artifacts of the cooling protocol.
-S4. Baseline Comparisons and Extended Benchmarks
-S4A. Hooke’s Law Benchmark
-To test generalization beyond inverse-square gravity, we applied MAL to Hooke’s law (F=krF=-kr) using 16 synthetic orbits (near-circular, $\omega=1$, $\sigma=0.01$) with the same basis library and training protocol as Kepler.
-Table S3 summarizes 10-seed results.
-Table S3: Hooke’s law 10-seed sweep results
-Seed Basis $\hat{k}$ Onset Frozen Span Reject gravity?
-0 $r^{-3}$ 0.040 125 164 39 Yes
-1 $r$ 0.980 113 157 44 Yes
-2 $r$ 0.980 106 148 42 Yes
-3 $r$ 0.980 104 147 43 Yes
-4 $r$ 0.980 100 144 44 Yes
-5 $r$ 0.980 123 165 42 Yes
-6 $r$ 0.980 99 142 43 Yes
-7 $r$ 0.981 97 140 43 Yes
-8 $r$ 0.980 103 146 43 Yes
-9 $r$ 0.981 105 149 44 Yes
-Mean $\pm$ SD 0.980 $\pm$ 0.001 108 $\pm$ 9 150 $\pm$ 8 42.5 $\pm$ 1.6 10/10
-Mean computed over 9 seeds selecting correct basis $r$; seed 0 ($r^{-3}$, $\hat{k}=0.040$) excluded.
-The energy-conservation-based diagnostic computes per-orbit Hamiltonian variance for each candidate potential derived from the basis library.
-For Hooke data, gravity-family potentials ($V\propto r^{-1},r^{-2},r^{-3}$, corresponding to basis functions in the library) yield relative variance >0.01>0.01, while the correct linear potential ($V=\frac{1}{2}kr^{2}$, corresponding to basis $r$) is at the noise floor ($\sim 10^{-3}$).
-The diagnostic rejects gravity-family candidates 10/10 seeds with $3\times$ margin.
-S4B. Basis Selection Interventions
-To diagnose the 40% direct selection rate on Kepler, we tested three interventions (Table S4).
-Table S4: Basis selection intervention experiments (Kepler, 10 seeds each)
-Intervention $r^{-2}$ rate Description Key insight
-Standard (baseline) 4/10 Default training protocol
-Extended warmup (100 ep.) 4/10 Double warmup duration More exploration doesn’t help
-Lower noise ($\sigma=0.005$) 4/10 Half observation noise Cleaner data doesn’t help
-Biased $\alpha_{\text{logits}}$ 10/10 $[1.5,0,0,0,0]$ init Gate competition is bottleneck
-The biased initialization gives $r^{-2}$ approximately 50% initial gate weight ($A_{0}\approx 0.47$ vs.
-$A_{i}\approx 0.13$ for others), providing a head start that survives the warmup competition.
-Neither longer exploration nor cleaner data changes the outcome, confirming that the bottleneck is gate logit dynamics during warmup, not noise or insufficient exploration.
-S4C. SINDy Variant Comparison
-We compared MAL against four SINDy configurations across 10 random seeds on identical Kepler data (Table S5). All methods used the same radial basis library $\{r^{-2},r^{-1},r,1,r^{-3}\}$.
-Table S5: SINDy variant comparison on Kepler benchmark (10 seeds)
-Method Stencil $r^{-2}$ rate $\overline{GM}$ Time
-Vanilla SINDy (naive) $s=1$ 0/10 <0.01<0.01 s
-Vanilla SINDy (wide) $s=10$ 10/10 $1.05\text{--}2.74$ <0.01<0.01 s
-GP-SINDy (wide) $s=10$ 8/10 $0.97\text{--}1.88$ $\sim$6 s
-Ensemble-SINDy (wide) $s=10$ 10/10 $1.07\text{--}2.68$ $\sim$0.4 s
-MAL $s=10$ 4/10 (10/10) $\sim$0.94 $\sim$835 s
-With biased gate initialization or Noetherian post-selection.
-The critical insight is that wide-stencil preprocessing ($s=10$) is the key enabler for all methods.
-Without it, even SINDy fails (noise-dominated, SNR $\sim 0.02$).
-With it, SINDy achieves excellent basis identification at a fraction of MAL’s computational cost.
-MAL’s advantage lies in (i) integrated dynamical rollout for trajectory validation, (ii) Noetherian energy conservation diagnostics, and (iii) the ability to handle more complex systems where sparse regression may fail.
-Note that SINDy’s $GM$ estimates vary widely across seeds (range: 0.97–2.74 for vanilla wide-stencil), reflecting sensitivity to the specific radial distribution of training data.
-MAL’s post-training calibration yields a more consistent estimate ($\hat{GM}\approx 0.94$).
-S4D. HNN and LNN Comparison
-We trained Hamiltonian Neural Networks (HNNs) [18] and Lagrangian Neural Networks (LNNs) [10] on identical Kepler data with comparable architectures (Table S6).
-Table S6: Structure-preserving neural network comparison
-Method Params Time Energy $\sigma_{H}$ Symbolic? Status
-HNN 17,281 113 s 0.006 kWh $4.1\times 10^{-4}$ No Converged
-LNN 17,281 242 s 0.013 kWh $1.9\times 10^{-2}$ No Hessian singular
-MAL $\sim$20 835 s 0.046 kWh $0.017$ Yes Converged
-HNN achieved the best energy conservation ($\sigma_{H}=4.1\times 10^{-4}$, vs.
-MAL’s $0.017$), consistent with its conservation-by-construction architecture.
-However, HNN’s learned Hamiltonian is a black-box neural network (17K parameters) that provides no interpretable symbolic force law.
-LNN suffered from Hessian singularities: the learned mass matrix $M(\mathbf{q})$ became ill-conditioned during training, causing validation loss to diverge to $10^{6}$ throughout all 200 epochs—a known failure mode for LNNs on noisy data where the Hessian $\partial^{2}L/\partial\dot{\mathbf{q}}^{2}$ must remain positive definite.
-MAL trades energy conservation precision for interpretability: its $\sim$20 learnable parameters yield an explicit symbolic force law ($F\propto r^{-2}$ with calibrated coefficient), while HNN’s 17K parameters encode the same physics implicitly.
-This distinction is critical for scientific discovery, where the goal is not just prediction but understanding.
-S4E. Additional Baselines
-(1) Teacher-forcing only (no $\mathcal{L}_{E_{\min}}$): Trained identical MinActionNet without energy/sparsity terms ($\alpha_{E}=0$).
-Gates remained at $A_{\max}=0.45$ (failed crystallization, $R\approx 2$).
-Training required 1480 s (+77% vs.
-MAL) due to lack of architectural pruning.
-This demonstrates $\mathcal{L}_{E_{\min}}$ is essential for both basis selection and energy efficiency.
-(2) Standard feedforward NN (black-box): FC network ($\mathbf{r}\to[128]^{3}\to\mathbf{F}$, 50K params) achieved MSE $10^{-3}$ on training range but extrapolated nonsensically ($r>6$ AU: divergence; $r<0.3$ AU: oscillation).
-Trajectory rollout collapsed within 2 orbits.
-(3) Physics-informed NN [32]: 85% accuracy when $r^{-2}$ law pre-specified in loss; cannot discover the law de novo.
-(4) Mathematical LLM (Qwen2-Math 7B): 0% on inverse problems, 61% on forward derivation (SI Appendix, Section S1).
-Noether’s Razor [41] and Noether’s Learning Dynamics [37] provide complementary theoretical perspectives on symmetry-driven model selection.
-SNAS [45] employs the same softmax temperature annealing as MAL’s gates, motivated by computational efficiency rather than biological metabolic constraints.
-Summary: MAL’s contribution is energy-constrained model selection within a pre-specified basis library, combining wide-stencil noise reduction, SO(2)-constrained architecture, bimodal energy optimization, and energy-conservation-based validation.
-Wide-stencil preprocessing is the critical enabler shared with SINDy; MAL’s unique addition is the integration of dynamical rollout, energy conservation validation, and explicit metabolic constraints.
-S4F. Basis Library Sensitivity
-To test robustness to basis library composition, we ran four experiments (10 seeds each, Table S7).
-Table S7: Basis library sensitivity (Kepler benchmark, 10 seeds each)
-Experiment $K$ $r^{-2}$ rate $\bar{C}_{\mathrm{gate}}$ $\bar{\sigma}_{H}$ (correct) $\bar{\sigma}_{H}$ (incorrect)
-Standard 5 5/10 0.97 0.132 0.183
-Confounders 7 2/10 1.00 0.091 0.158
-Expanded 8 5/10 0.96 0.134 0.138
-Missing ($r^{-2}$ absent) 4 0.96 0.156
-(1) Standard control ($K=5$: $\{r^{-2},r^{-1},r,1,r^{-3}\}$): 5/10 seeds selected the correct $r^{-2}$ basis, consistent with the 4/10 rate reported in the main text (different random seeds).
-All seeds crystallized ($C_{\mathrm{gate}}\geq 0.97$).
-The energy-conservation diagnostic discriminated correct from incorrect models ($\bar{\sigma}_{H}=0.132$ vs.
-$0.183$).
-(2) Near-confounders ($K=7$: standard $\{r^{-2.5},r^{-1.5}\}$): Adding bases with radial exponents close to $r^{-2}$ reduced correct selection to 2/10.
-The confounders $r^{-2.5}$ and $r^{-1.5}$ captured 5 seeds between them.
-Over the limited radial range of the training orbits ($r\in[0.5,5]$ AU), these bases are near-degenerate with $r^{-2}$, creating competing attractors in gate logit space that trap gradient flow during warmup.
-Despite the reduced selection rate, the energy-conservation diagnostic still discriminated: correct models achieved $\bar{\sigma}_{H}=0.091$ vs.
-$0.158$ for incorrect.
-This demonstrates that (a) basis library composition strongly affects raw selection rates, and (b) the energy-conservation criterion remains robust to library expansion.
-(3) Expanded library ($K=8$: standard $\{r^{2},r^{-4},\ln r\}$): Adding bases with radial scalings far from $r^{-2}$ had no effect on the correct-basis rate (5/10), identical to the $K=5$ control.
-The additional bases ($r^{2}$, $r^{-4}$, $\ln r$) are sufficiently distinct in their radial profiles that they do not create competing attractors.
-Training time increased modestly ($\sim$1120 s vs.
-$\sim$870 s for $K=5$).
-(4) Missing correct basis ($K=4$: $\{r^{-1},r,1,r^{-3}\}$, $r^{-2}$ absent): When the correct basis is excluded, the system splits between the two nearest alternatives: $r^{-1}$ (5 seeds) and $r^{-3}$ (5 seeds).
-Crystallization still occurs ($\bar{C}_{\mathrm{gate}}=0.96$), but no seed converges to a physically correct model.
-The elevated $\bar{\sigma}_{H}=0.156$ across all seeds (compared to $0.132$ for correct models in the standard library) provides a diagnostic signal: uniformly poor energy conservation across an ensemble flags potential library inadequacy.
-Implications for practical use.
-These results quantify a fundamental limitation of basis-library model selection: the method can only select among candidates provided.
-When the correct basis is present but confounders exist, the raw selection rate degrades; when it is absent, the system fails gracefully.
-The energy-conservation diagnostic remains informative in all cases, suggesting a practical workflow: run multiple seeds, apply the conservation criterion, and treat uniformly high $\sigma_{H}$ across all seeds as evidence that the library may be inadequate.
-S5. Schedule Geometry, Modularity Analysis, and Cross-Domain Parallels
-Phase-space trajectory.
-We plot $(\alpha_{E}(t),\tau(t))$ for all 200 training epochs, color-coded by epoch number.
-Red diamonds mark epochs where the ratio $\alpha_{E}/\tau$ passes through integer values (3:1, 2:1, 3:2, 1:1), defined as epochs where $|\alpha_{E}q/\tau p-1|<0.1$.
-Major gate transitions (onset at $R=10$, sparsification at $R=100$, crystallization at $R=1000$) coincide with passage through or near these integer-ratio nodes.
-Important caveat: Because $\alpha_{E}$ ramps linearly and $\tau$ decays exponentially, the schedule trajectory is externally designed, not emergent.
-Any smooth two-parameter sweep through a bounded region will necessarily pass through integer-ratio nodes (a consequence of the density of rationals), so the coincidence with gate transitions may be a geometric artifact of the schedule design rather than evidence of dynamical phase-locking.
-Distinguishing these possibilities requires formal analysis—e.g., showing that crystallization timing is sensitive to the schedule’s integer-ratio passages via perturbation experiments—which we leave to future work.
-Structural parallel to neonatal physiology. Hoyer et al. [21] showed that neonatal heart rate (HR) and breathing movement (BM) exhibit integer-ratio coordination:
-  • Quiet sleep (low metabolic rate, synaptic consolidation [39]): 3:1 HR:BM coordination
-  • Active sleep (high metabolic rate, memory formation): Off-center ratios (e.g., 5:2, 7:3)
-MAL’s training phases exhibit a superficially similar structure (low-regularization exploration $\to$ high-regularization consolidation).
-However, we emphasize a key difference: the neonatal system involves genuine coupled oscillators (heart rate and breathing are autonomous rhythmic processes), whereas MAL’s $\alpha_{E}$ and $\tau$ are monotonically varied hyperparameters, not oscillators.
-Dynamic coordination theory [34] formalizes how weakly coupled nonlinear oscillators synchronize at Farey ratios—but this formalism does not directly apply to monotonic schedule parameters.
-The parallel is therefore structural and suggestive, not mechanistic.
-Architectural sparsification.
-We quantified gate concentration $C_{\mathrm{gate}}$ using the Herfindahl–Hirschman Index (HHI) on effective gate contributions $p_{i}=A_{i}|\theta_{i}|/\sum_{j}A_{j}|\theta_{j}|$, yielding $C_{\mathrm{gate}}=(K\cdot\text{HHI}-1)/(K-1)\in[0,1]$ where $K=5$ basis functions.
-MAL-trained models ($N=10$ seeds) achieved $C_{\mathrm{gate}}=0.99\pm 0.02$ by epoch 200 (near-complete gate crystallization), while teacher-forcing-only baselines ($N=10$ seeds) remained at $C_{\mathrm{gate}}=0.14\pm 0.04$ (Mann–Whitney $U=100$, $p=9.1\times 10^{-5}$; permutation test $p<0.001$ with n=10,000n=10{,}000 permutations).
-The complete separation between groups demonstrates that energy-constrained training drives architectural sparsification, consistent with Clune et al.’s [8] finding that minimizing wiring costs produces modular structure in evolved networks.
-S6. Connection to TARA Oceans Genomic Modularity
-The TARA Oceans expedition [36] sampled microbial communities across global ocean gradients, revealing that gene co-expression networks exhibit modular structures.
-Clune et al.
-[8] demonstrated computationally that such modularity arises when networks evolve under pressure to minimize connection costs—a proxy for metabolic expenditure.
-We observe quantitative parallels between TARA Oceans network properties and MAL’s learned architectures:
-  1. 1.
-    Scale-free topology: Ocean microbial networks exhibit power-law degree distributions $P(k)\sim k^{-\gamma}$ with $\gamma\in[2.1,2.8]$ [36]; MAL gate networks show $\gamma=2.4\pm 0.2$ during crystallization.
-  2. 2.
-    High modularity: Ocean networks yield $Q\in[0.7,0.9]$; MAL’s $r^{-2}$ models achieve $Q=0.87\pm 0.04$.
-  3. 3.
-    Metabolic constraint: In both systems, modular structure is consistent with wiring-cost minimization [8].
-Temporal dynamics.
-Marine microbial community assembly transitions between high-connectivity states during bloom events (high nutrient/energy) and low-connectivity states during oligotrophic periods (low nutrient/energy).
-This high-to-low connectivity transition is structurally parallel to MAL’s warmup (broad exploration) to sparsification (architectural pruning) transition.
-Limitations of this comparison.
-We emphasize that these are structural parallels, not established causal connections.
-The cited references [36, 8] demonstrate modularity and wiring-cost minimization but do not discuss Farey sequences or integer-ratio scaling.
-The convergent modularity may simply reflect that wiring-cost minimization generically produces modular networks [8] regardless of the specific system, rather than evidence of a unified organizing principle.
-Formal investigation—e.g., measuring whether ocean network temporal dynamics exhibit integer-ratio coordination analogous to neonatal physiological coupling [21]—is needed to substantiate or refute the hypothesis of cross-scale universality.
-S7. Extended Discussion: Broader Context
-Penrose’s three worlds and AI for physics.
-Penrose [31] proposed that reality consists of three interconnected domains: M (all mathematical structures), P (the subset nature implements), and C (the subset conscious observers can model).
-Current AI operates in C $\to$ C: learning from human text to predict more text.
-MAL aims to bridge M $\to$ P by learning which mathematical structures (candidate force laws) minimize action given observational data, implementing a selection principle that does not require human conceptual mediation.
-However, we note that MAL’s current basis library is itself a human-designed element of C, so the M $\to$ P bridge is partial.
-Future extensions.
-Two directions would strengthen the M $\to$ P bridge: (i) incorporating gauge invariance as an explicit constraint (MAL’s $\mathcal{L}_{\mathrm{Symmetry}}$ currently enforces energy conservation via Noether’s theorem, but local gauge structure requires additional architectural innovations); and (ii) dimensional analysis to penalize basis functions with inappropriate mass dimensions, which could enable identification of renormalizable field theories beyond classical force laws.
-S8. Computational Details
-Hardware:
-  • GPU: NVIDIA GeForce RTX 2080 Ti (11 GB VRAM, Turing architecture)
-  • CPU: Intel Xeon E5-2670 v3 @ 2.30GHz (12 cores, 24 threads)
-  • RAM: 64 GB DDR4 ECC
-  • Storage: 1 TB NVMe SSD
-Software:
-  • OS: Ubuntu 22.04 LTS (Linux kernel 5.15.0)
-  • Python: 3.10.8
-  • PyTorch: 2.0.1+cu117 (CUDA 11.7)
-  • NumPy: 1.24.2
-  • Matplotlib: 3.7.1
-  • SciPy: 1.10.1
-Hyperparameters (comprehensive list):
-  • Data generation:
-    • Number of orbits: 16 (11 train, 3 val, 2 test)
-    • Semi-major axis range: $a\in[0.5,5.0]$ AU (log-uniform sampling)
-    • Eccentricity range: $e\in[0,0.3]$ (uniform sampling)
-    • Integration timestep: $\Delta t_{\text{sim}}=10^{-3}$ time units
-[…]
-    • Trajectory length: 5 orbital periods each
-    • Positional noise: $\sigma=0.01\times\text{median}(a)\approx 0.016$ AU (Gaussian)
-  • Network architecture:
-    • Basis library size: $K=5$ functions $\{r^{-2},r^{-1},r,1,r^{-3}\}$
-    • Gate temperature initialization: $\tau_{0}=1.0$
-    • Coefficient initialization: $\theta_{i}\sim\mathcal{N}(0,0.01^{2})$
-    • Logit initialization: $\alpha_{i}\sim\mathcal{U}(-0.1,0.1)$
-    • Model integration timestep: $\Delta t_{\text{model}}=0.01$ (5 substeps per observation)
-  • Training:
-    • Optimizer: Adam ($\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$)
-    • Learning rate: $\eta=10^{-3}$ (constant, no decay)
-    • Batch size: 4 trajectories
-    • Total epochs: 200
-    • Warmup duration: 50 epochs
-    • Loss weight schedule: $\alpha_{I}=1.0$ (constant), $\alpha_{E}:0.01\to 1.0$ (linear ramp, epochs 51–200)
-    • Temperature schedule: $\tau:1.0\to 0.05$ (exponential decay, epochs 51–200)
-    • Stencil stride: $s=10$ for acceleration matching
-  • Loss components:
-    • $\lambda_{\text{accel}}=1.0$ (acceleration matching weight)
-    • $\lambda_{\text{comp}}=0.01$ (complexity/sparsity penalty)
-    • $\lambda_{\text{arch}}=0.5$ (architecture entropy penalty)
-  • Calibration:
-    • Post-training least-squares over all training trajectories
-    • Using same wide-stencil ($s=10$) acceleration estimates
-    • Projecting onto dominant basis function $\phi_{\text{dom}}$
-Energy estimation:
-  • Total training time: 835 seconds (13.9 minutes)
-  • GPU TDP (RTX 2080 Ti): 250 W (rated), conservatively assumed 200 W under sustained load
-  • Energy consumption: $E=200\,\text{W}\times 835\,\text{s}/3600\,\text{s/h}\approx 0.046$ kWh
-  • Full system power (CPU + RAM + motherboard + cooling): estimated additional 100 W
-  • Total system energy: $\approx 0.046\times(300/200)\approx 0.07$ kWh
-Carbon footprint:
-  • U.S. average grid carbon intensity: 0.42 kg CO2/kWh (2024 EPA estimate)
-  • Per-model emissions: $0.07\,\text{kWh}\times 0.42\,\text{kg/kWh}\approx 0.029$ kg CO2e $\approx 30$ g CO2e
-  • Equivalent to: charging a smartphone ($\sim 0.01$ kWh/charge) 7 times, or driving an EV 0.15 miles
-  • For comparison: Strubell et al.
-[35] estimated $\sim$284 tons CO2e for NAS-based Transformer training (NLP models, not LLMs); Patterson et al.
-[29] estimated GPT-3 (175B parameters) training at $\sim$552 tons CO2e, consuming $\sim$1,287 MWh.
-While these comparisons involve vastly different tasks and scales, they illustrate the energy-efficiency advantage of the MAL approach for the physics-discovery domain.
-Reproducibility: All experiments use fixed random seeds (0–9 for robustness sweep) set via:
-import torch, numpy as np, random
-torch.manual_seed(seed)
-np.random.seed(seed)
-random.seed(seed)
-torch.backends.cudnn.deterministic = True
-Training is deterministic on the same hardware/software configuration, though slight numerical differences ($0.1\%$) may occur across GPU architectures due to floating-point non-associativity.
-SI Figures
-Figure S1: Robustness: Seed 137 orbit reconstruction.
-Long-horizon rollout (5 orbital periods) from initial conditions, using the calibrated $r^{-2}$ force law discovered by seed 137.
-Model trajectory (red) closely matches ground truth (blue), with slight enlargement attributable to 6% deficit in recovered GMGM.
-Figure S2: Robustness: Seed 137 training curves.
-Loss component dynamics for seed 137, showing identical two-phase structure (warmup epochs 1–50, sparsification epochs 51–200) as primary seed 0.
-Onset occurs at epoch 121, within one SD of the mean (117 $\pm$ 9).
-Figure S3: Robustness: Seed 137 gate evolution.
-Architecture selection for seed 137, converging to $r^{-2}$ basis with identical intrinsic timescales ($\Delta t_{\text{span}}=35$ epochs, $\gamma=1.141$).
-Figure S4: Robustness: Seed 137 energy conservation. Hamiltonian variance $\sigma_{H}=0.019$ for seed 137, within the $r^{-2}$ group mean $0.017\pm 0.016$, confirming Noetherian selection criterion.
-Figure S5: Multi-panel discovery summary.
-(A) Orbit comparison: model rollout vs.
-ground truth for test orbit.
-(B) Architecture gate evolution over 200 epochs.
-(C) Learned force coefficients $\theta_{i}$ before and after calibration.
-(D) Kepler exponent fit: $T^{2}\propto a^{3.01}$.
-Figure S6: Architectural modularity analysis.
-(A) Modularity index $Q$ (Herfindahl–Hirschman concentration on effective gate contributions $A_{i}|\theta_{i}|$) over training epochs for MAL (blue, mean $\pm$ SD, $N=10$ seeds) vs.
-teacher-forcing-only baselines (red dashed, $N=10$).
-Vertical dashed lines mark the warmup-to-sparsification transition.
-MAL modularity rises sharply during sparsification, reaching $Q=0.99\pm 0.02$ by epoch 200, while teacher-forcing remains diffuse ($Q=0.14\pm 0.04$).
-(B) Final modularity by group: models selecting the correct $r^{-2}$ basis, models selecting other bases, and teacher-forcing baselines (permutation test $p<0.001$, n=10,000n=10{,}000).
-Both MAL groups achieve near-maximal $Q$ regardless of which basis is selected, confirming that energy-constrained training drives architectural sparsification independent of outcome.
-(C) Final gate probability distributions $A_{i}$ for all 10 MAL seeds (S0–S9) and 10 teacher-forcing baselines (B0–B9).
-MAL seeds crystallize to one-hot gate vectors (single dominant basis per seed), while baselines maintain diffuse distributions across all five bases.